### v1

# From Fibers to Cells: Fourier-Based Registration Enables Virtual Cresyl Violet Staining From 3D Polarized Light Imaging

Abstract
Comprehensive assessment of the various aspects of the brain’s microstructure requires the use of complementary imaging techniques. This includes measuring the spatial distribution of cell bodies (cytoarchitecture) and nerve fibers (myeloarchitecture). The gold standard for cytoarchitectonic analysis is light microscopic imaging of cell-body stained tissue sections. To reveal the 3D orientations of nerve fibers, 3D Polarized Light Imaging (3D-PLI) has been introduced as a reliable technique providing a resolution in the micrometer range while allowing processing of series of complete brain sections. 3D-PLI acquisition is label-free and allows subsequent staining of sections after 3D-PLI measurement. By post-staining for cell bodies, a direct link between fiber- and cytoarchitecture can potentially be established within one and the same section. However, inevitable distortions introduced during the staining process make a costly nonlinear and cross-modal registration necessary in order to study the detailed relationships between cells and fibers in the images. In addition, the complexity of processing histological sections for post-staining only allows for a limited number of samples. In this work, we take advantage of deep learning methods for image-to-image translation to generate a virtual staining of 3D-PLI that is spatially aligned at the cellular level. We use a supervised setting, building on a unique dataset of brain sections, to which Cresyl violet staining has been applied after 3D-PLI measurement. To ensure a high correspondence between both modalities, we address the misalignment of training data using Fourier-based registration methods. In this way, registration can be efficiently calculated during training for local image patches of target and predicted staining. We demonstrate that the proposed method enables prediction of a Cresyl violet staining from 3D-PLI, matching individual cell instances.
Keywords deep learning, virtual staining, fiber architecture, cytoarchitecture, polarized light imaging, Cresyl violet, vervet monkey brain
1 Introduction
To understand the organizational principles of the brain, complementary imaging techniques are used to highlight different aspects of brain architecture. Two important aspects of the microstructural organization are fiber- and cytoarchitecture Nieuwenhuys, (2013); Amunts and Zilles, (2015). While cytoarchitecture encompasses the spatial distribution and shape of cell bodies in the cerebral cortex and subcortical nuclei, fiber architecture refers to the course and composition of nerve fibers. However, cytoarchitecture and fiber-architecture are usually studied using different staining protocols, applied in different sections. Only a few protocols are available to combine cyto- and fiber staining in a single protocol, e.g., Luxol fast blue Klüver and Barrera, (1953), Bielschowsky Bielschowsky, (1904) or the triple staining by Novotny Novotny and and Novotny, (1977) and Nowotny (2009). While they allow visualizing cell bodies and fibers in one and the same section, they do not allow the tracing of axons and fiber bundles over long distances.
3D-Polarized Light Imaging (3D-PLI) addresses this limitation. It is a microscopic imaging technique for evaluating the three-dimensional orientation of myelinated nerve fibers in entire, unstained histological brain sections Axer and Amunts, (2022); Axer et al., 2011a ; Axer et al., 2011b . The technique can achieve an in-plane resolution of 1.3 µm, capturing structures at the level of individual fibers and small fiber bundles. 3D-PLI has been used to gain insights into the architecture of nerve fibers in different brain regions, such as the human hippocampus Zeineh et al., (2017), the sagittal stratum Caspers et al., (2022) and the vervet monkey visual system Takemura et al., (2020). In addition, 3D-PLI has been used to validate fiber tractography algorithms and the interpretation of DW-MRI Caspers and Axer, (2019). 3D-PLI potentially allows joint imaging of fiber tracts and neuronal cell bodies Zeineh et al., (2017) due to scattering effects affecting light transmittance Menzel et al., (2020). However, this possibility has not yet been validated.
Cytoarchitecture can be studied in histological sections of postmortem brains with Cresyl violet. The staining provides contrast due to staining of the rough endoplasmic reticulum. This allows to study cell shape, density and distribution, which vary between brain regions. Due to its high spatial resolution, microscopic analysis of histological sections is considered the gold standard to verify structural parcellations Amunts and Zilles, (2015). Recent advances in high-throughput scanning, data processing algorithms, and computational capacities have enabled the creation of 3D human brain atlases based on cytoarchitecture, such as BigBrain Amunts et al., (2013), the Allen Adult Human Brain Atlas (AAHA, Ding et al., 2016), and the Julich Brain probabilistic atlas Amunts et al., (2020).
Since 3D-PLI relies solely on optical properties of the tissue, it is label-free and can be combined with a staining of the same tissue after its measurement. Post-staining, e.g. with Cresyl violet, enables a complementary visualization of neuronal cell bodies, potentially establishing a direct link between cytoarchitecture with 3D fiber-architecture. This requires, however, a complex histological processing, which limits the number of available samples, increases the risk of tissue damage and may lead to deformations of the section. To correct for deformation and artifacts in the images of the two modalities requires a nonlinear registration step. However, 3D-PLI and Cresyl violet stained tissue share only a few automatically identifiable cross-modal registration landmarks, such as blood vessels or morphological landmarks. I.e., post-staining of sections imaged with 3D-PLI is feasible, but technically challenging, and it hardly scales up.
Figure 1: The proposed virtual staining workflow. (A) Preprocessing of 3D-PLI sections that were post-stained with Cresyl violet. As paired training data, regions of interest (ROIs) are manually cropped (red boxes) and affine registered using large blood vessels as landmarks (green marker). Background pixels in train sections are masked, and retardation values are scaled using gamma correction for visualization purposes. (B) Training of a U-Net model using patches, extracted from same random locations (yellow boxes) in 3D-PLI modalities direction $\varphi$, retardation $\sin\delta$, transmittance $T_{T}$ and the Cresyl violet staining. 3D-PLI patches are used as input to the model to predict a virtual Cresyl violet staining. The Cresyl violet patch acts as target and is rigidly aligned with the prediction during the training procedure. The alignment is performed by our proposed online registration head using Fourier-based correlation of pixels. A loss $\mathcal{L}$ is computed between aligned target and prediction. (C) Inference using the trained U-Net model to virtually stain unseen sections or ROIs. Inputs are divided into overlapping tiles, which are processed independently by the U-Net model. The predictions are then stitched back together to form the complete virtual staining.
Therefore, we aim to train a deep neural network model to perform image-to-image translation from 3D-PLI to a Cresyl violet staining. Such an approach is often denoted as virtual histological staining, which refers to computational methods that generate color-coded images of biological tissue without the need for traditional staining techniques. The methods instead utilize optical properties of the tissue, such as birefringence, autofluorescence, scattering, or absorption to create images that emulate the appearance of stained tissue. A virtual Cresyl violet staining spatially aligned with 3D-PLI would allow an application of established tools for cytoarchitectonic analysis, such as automatic cell instance segmentation Upschulte et al., (2022), for 3D-PLI measurements. Furthermore, it would provide detailed registration landmarks for cross-modal registration Ounkomol et al., (2018), thereby offering the opportunity to perform joint acquisition of aligned fiber and cytoarchitecture at a larger scale than possible today.
One of the earliest applications of label-free imaging for visual staining was Quantitative Phase Imaging (QPI, Curl et al., 2004). QPI measures the phase shift of light passing through a sample, producing high-resolution images that reveal the optical properties of the tissue. It was used to generate images of collagen fibers, red blood cells, and other tissue structures without staining Curl et al., (2004); Park et al., (2018). Later, machine-learning algorithms, especially generative models, have been trained to recognize and virtually stain different tissue structures in unstained images by performing image-to-image translation. They have successfully generated color-coded images of tissue that replicate the appearance of histological stainings, such as a virtual hematoxylin and eosin (H&E), Masson’s trichrome, and Jones’ stain from QPI of label-free tissue Rivenson et al., (2019), a transformation of H&E stained tissue into Masson’s trichrome, periodic acid-Schiff (PAS), or Jones’ stain de Haan et al., (2021); Yang et al., (2022). However, these stains are not very good at distinguishing the different components of the nervous tissue.
Machine learning algorithms were also used to predict fluorescence-labeled images from transmitted-light z-stacks Ounkomol et al., (2018); Christiansen et al., (2018); Cross-Zamirski et al., (2022) or 3D fluorescence structures and a FluoroMyelin stain from bright-field and polarization images of brain slices Guo et al., (2020). The methods typically use cross-entropy Christiansen et al., (2018), mean absolute (L1) loss Guo et al., (2020), or style-related losses such as conditional generative adversarial network (GAN) loss Rivenson et al., (2019); de Haan et al., (2021); Cross-Zamirski et al., (2022). While including a GAN objective encourages prediction of realistic-looking images, it has no clear mechanism to preserve content when conditioned on a particular input image, and thus may introduce artificial structures Cohen et al., (2018). A combination with a pixel-wise reconstruction loss (e.g. L1 loss) mitigates this problem of GAN training and improves accuracy of predictions Isola et al., (2017).
Since methods using paired training data for supervised image-to-image translation typically produce more accurate predictions than unpaired methods Zhu et al., (2017), a pixel-accurate alignment of training data is necessary. This requires virtual staining methods to either perform a costly registration step or directly acquire paired images. A paired acquisition with 3D-PLI, however, is not feasible and a lack of structural overlap, such as a sufficient number of visible cell instances between the investigated modalities, makes pixel-accurate registration challenging. Therefore, to alleviate the need for perfectly paired training data, we propose a supervised learning objective performing local online registration of training pairs combined with a translational-invariant style comparison. This allows us to train the model on imperfectly registered image pairs with strong content preservation as in paired image-to-image translation, while enabling realistic prediction of subtle structures like cell bodies (Fig. 1). The main contributions are as follows:
  • We apply the matching of Gram matrix representations as a texture sensitive style loss for the virtual staining, as previously used for texture synthesis Gatys et al., (2015). Since the computation of Gram matrices is translation invariant, it allows a direct comparison of image statistics between coarsely aligned training examples. It therefore improves the accuracy of predicted cell instances over commonly used GAN style loss.
  • An online registration head for improving registration accuracy of local image pairs after pre-alignment of larger tissue sections during training. We consider Fourier-based registration methods, which can be computed efficiently in real-time on modern GPU hardware.
  • An equivariance loss to improve the accuracy of cell instance predictions by addressing the inherent agnosticism of loss computation to constant displacements through online registration.
2 Materials and Methods
2.1 Microscopic imaging of histological brain sections
We demonstrate the proposed virtual staining approach on a set of brain sections for which Cresyl violet staining has been performed after 3D-PLI measurement. The brain sample for this study was obtained from a healthy 2.4-year-old adult male vervet monkey (Wake Forest-ID 1818, Axer et al., 2020; Takemura et al., 2020; Menzel et al., 2022) in accordance with the Wake Forest Institutional Animal Care and Use Committee (IACUC #A11-219) and conforming the AVMA Guidelines for the Euthanasia of Animals. To obtain an undistorted volumetric reference, a T2 weighted MRI was acquired in-vivo one day prior to sacrifice. The brain was removed from the skull within 24 hours after flushing with phosphate-buffered saline and perfusion fixated with 4% buffered paraformaldehyde. It was stored for several weeks at -70°C in 20% glycerin solution for cryoprotection, and then sliced coronally with 60 µm section thickness using a large-scale cryostat microtome (Polycut CM 3500, Leica Microsystems, Germany). Blockface images of the frozen tissue block were taken with a CCD camera before cutting each brain section. The images were reconstructed into a 3D blockface volume to provide an undistorted reference for section realignment.
2.1.1 Image acquisition
Figure 2: Data modalities and registration challenges for training section 544. (A-C) 3D-PLI parameter maps: Transmittance $I_{T}$, retardation $\sin\delta$ and fiber orientation in HSV color space (hue: direction $\varphi$; saturation, brightness: retardation $\sin\delta$). Background pixels are masked for visualization purposes only. (D) Affine registered Cresyl violet staining. The pial surface of the 3D-PLI acquisition is shown as a contour plot in D for reference. Between both data acquisitions remains a nonlinear misalignment that cannot be resolved by a global affine transformation. At a local scale, the remaining misalignment is approximately linear. Yellow arrows indicate blood vessels that can be used as mutual registration landmarks for coarse alignment.
For 3D-PLI acquisition, brain sections were scanned using a polarizing microscope (LMP-1, Taorad, Germany) with 1.3 µm resolution Axer et al., 2011b ; Axer and Amunts, (2022). The focus level of the LMP-1 was manually adjusted to the center of the tissue for each section. Inside the LMP-1 microscope, sections were placed on a specimen stage between a rotating linear and a circular polarizer on top of an incoherent light source with a wavelength of 550 $\pm$ 5 nm. Images were taken by a CCD camera for nine equidistant rotation angles $\rho$ of the rotating linear polarizer, covering 180° of rotation. At each pixel, the measured intensity of the images followed a sinusoidal profile as
$I_{\rho}=\frac{I_{T}}{2}\left(1+\sin(2\rho-2\varphi)\sin\delta\right).$ (1)
Using harmonic Fourier analysis, parameter maps of transmittance ($I_{T}$), retardation ($\sin\delta$), and direction ($\varphi$) were obtained from the measurements, with a resolution of approximately 34,000 $\times$ 44,000 pixels per section, revealing their fine-grained nerve fiber architecture (Fig. 2A-C). After 3D-PLI acquisition, brain sections were washed, fixed and stained for cell bodies with Cresyl violet Nissl staining to reveal their cellular architecture. They were scanned using flatbed scanners at 1 µm in-plane resolution and saved as RGB color pictures with eight bit color depth, storing pixel values from 0 to 255 (Fig. 2D).
2.1.2 Tissue shrinkage correction
Deformations induced by histological processing include a volumetric shrinkage of brain tissue Amunts et al., (2005). To correct this effect, the extent of shrinkage is usually estimated from the ratio between the histological processed and true brain volume represented by the fresh weight of the whole brain Amunts et al., (2005, 2013) or an MRI reference Wagstyl et al., (2020). To estimate the shrinkage of the 3D-PLI acquisition, a linear registration of the 3D reconstructed blockface volume and the postmortem MRI of the same brain was performed. The determinant of the transformation matrix showed a global volume change below 1%, which can be neglected. In a second step, we calculated 2D shrinkage factors for each brain section considered in this work as the quotient between the area occupied by tissue in the 3D-PLI measurement and its corresponding blockface image. For test section #559, we estimated a 2D shrinkage factor of 0.97 meaning a slight expansion of the 3D-PLI measurement relative to its original area within the MRI. This observation is consistent throughout all sections with 2D shrinkage factors between 0.95 and 0.99. We report in-plane cell sizes measured with 3D-PLI multiplied by an individual 2D shrinkage factor per section, assuming a uniform area change of cells and surrounding tissue.
2.2 Initial cross-modality alignment
After the subsequent processing of brain tissue, Cresyl violet images exhibit a deformation relative to the 3D-PLI acquisition. To align both modalities, an initial affine registration of whole brain sections is performed by manual identification of large blood vessels as landmarks visible in both modalities.
Performing the initial affine registration reveals remaining nonlinear deformations as shown in Fig. 2D. Since nonlinear deformations typically have low spatial frequencies, causing smooth, large-scale distortions, we expected near-linear deformations at smaller scales. Therefore, performing an additional more local linear registration would lead to a better fit. We subsequently crop square regions of interest (ROIs) with a size of 4,096 pixels ($\sim$5.3 mm) and without visible artifacts, covering distinct cellular architectures across the whole coronal plane. For all ROIs we perform additional affine registration and make sure that transformed landmarks have a maximum distance of 70 pixels (91 µm) from their matches. All ROIs are warped and resampled to 1.3 µm using linear interpolation to match the coordinate space of 3D-PLI. While this results in a loss of precision relative to the original resolution of Cresyl violet scans of 1 µm, matching the resolutions of both modalities facilitates subsequent processing and analysis steps.
2.3 Fourier-based online registration of image patches
To correct the remaining misalignment of 3D-PLI and Cresyl violet after affine registration of ROIs at a finer local scale, we introduce an online registration head that performs cross-modality alignment during training based on model predictions of small image patches. We assume that once the style transfer model has learned to reconstruct microscopic landmarks (e.g., individual cells or small blood vessels), such online registration will promote the learning of additional landmarks until the training can use pixel-aligned training examples.
The registration method performed during training needs to be computationally efficient, since training will require numerous registration iterations. Conventional feature-based image registration methods are accurate and can model nonlinear deformations but are computationally expensive and sensitive to image degradation. As we assume deformations to be approximately linear at a local scale, we take advantage of Fourier-based image correlations Tong et al., (2019), which can efficiently recover a translation between images in the frequency domain.
2.3.1 Translational shift
Fourier-based image correlation methods are able to retrieve a translational shift $(\Delta u,\Delta v)$ between image functions $f(u,v)$ and $g(u,v)$ defined for integer pixel coordinates $(u,v)$, such that $f(u,v)=g(u+\Delta u,v+\Delta v)$. Both functions $f$ and $g$ represent images of equal height $H$ and width $W$ and are for now assumed to repeat periodically with a periodicity of $H$ and $W$, respectively.
A common approach to retrieve the translational shift between $f$ and $g$ is to use circular cross-correlation Tong et al., (2019), which can be efficiently computed in the frequency domain as
$\begin{split}\text{CC°}[a,b]&=(f\star g)[a,b]\\ &=\sum_{u,v}f(u-a,v-b)g(u,v)\\ &=\left(\mathcal{F}^{-1}\left\{\overline{\mathcal{F}\{f\}}\mathcal{F}\{g\}\right\}\right)[a,b],\end{split}$ (2)
for all integer shifts $[a,b]$, where $\mathcal{F}$ denotes the Fourier transformation, $\mathcal{F}^{-1}$ its inverse, $\overline{\mathcal{F}\{f\}}$ its complex-conjugate Fourier coefficients, and where we sum over all pixel coordinates $(u,v)$. The translational shift can subsequently be recovered by the location of the maximum value in CC° as
$(\Delta u,\Delta v)=\operatorname*{arg\,max}_{(a,b)}\text{CC°}[a,b].$ (3)
Eq. 2 can be extended to the mean over squared distances between pixel values as
$\begin{split}\text{MSE°}[a,b]&=\frac{1}{HW}\sum_{u,v}(f(u-a,v-b)-g(u,v))^{2}\\ &=\frac{-2(f\star g)[a,b]+\sum_{u,v}(f^{2}(u,v)+g^{2}(u,v))}{HW}.\end{split}$ (4)
The translational shift $(\Delta u,\Delta v)$ can be recovered analog to Eq. 3 by computing the $\operatorname*{arg\,min}$. For periodic image functions $f$ and $g$, solutions of CC° and MSE° are identical as the sum over squared functions $f^{2}$ and $g^{2}$ is constant Fienup, (1997).
Since histological images are not periodic, zero padding is applied to fill both images up to a shape of $(H_{f}+H_{g}-1,W_{f}+W_{g}-1)$, in order to break periodicity and allow processing of images with different heights $H_{f}$, $H_{g}$ and widths $W_{f}$, $W_{g}$. The zero-padded images are denoted by new image functions $f_{0}$ and $g_{0}$. Furthermore, additional masks $M_{f}$ and $M_{g}$ are introduced, which have a value of one at all pixel coordinates within original height and width and zero elsewhere. We reformulate Eq. 4 to a non-circular form as
$\text{MSE}[a,b]=\left(\frac{(f_{0}^{2}\star M_{g})-2(f_{0}\star g_{0})+(M_{f}\star g_{0}^{2})}{M_{f}\star M_{g}}\right)[a,b],$ (5)
which can be efficiently computed by multiple applications of Eq. 2. Here, the cross-correlations of $f_{0}^{2}$ and $g_{0}^{2}$ with masks $M_{g}$ and $M_{f}$, respectively, ensure that pixel values of the original unpadded images are not compared with zero padding values. Furthermore, the score is divided by the correlation between $M_{f}$ and $M_{g}$ to account for the number of overlapping pixels between original unpadded images. We apply the same division by $M_{f}\star M_{g}$ also for CC° to retrieve a non-circular variant called CC. While the solutions for circular CC° and MSE° are identical, the solutions for non-circular CC and MSE differ, resulting in distinct registration metrics.
2.3.2 Rotation and scale
While relative scale and rotation between images can also be retrieved in the frequency domain Sheng, (1989); Reddy and Chatterji, (1996), we expect only small relative rotation angles and minor scale variations due to the initial affine registration by matching blood vessels. Therefore, we leverage the parallel processing capability of GPUs to perform an exhaustive search over a fixed set of rotation angles without scaling adjustments. We calculate registration metrics for every translational shift and rotation and select the rotation angle and shift combination that yields a global optimum.
Figure 3: Illustration of the proposed virtual staining approach. Patches of 3D-PLI parameter maps transmittance $I_{T}$, retardation $\sin\delta$ (scaled using gamma correction for visualization), and direction $\varphi$ are used as input to a 2D convolutional U-Net model as generator to predict a virtual Cresyl violet staining. An online registration head estimates a rigid transformation $R$ between a coarsely aligned Cresyl violet target patch and the prediction via Fourier-based registration. Transformation $R$ is used to align target and prediction at the patch level. We calculate three distinct loss components: $\mathcal{L}_{R}$, $\mathcal{L}_{S}$ and $\mathcal{L}_{E}$. Reconstruction loss $\mathcal{L}_{R}$ performs a pixel-wise comparison between prediction and aligned target. Style loss $\mathcal{L}_{S}$ compares feature maps of a VGG network encoder using Gram matrices to mimic the style of the target image. Equivariance loss $\mathcal{L}_{E}$ applies the same U-Net model a second time to a rotated version of the input by rotation $\Omega$. The output is compared with the prediction rotated by same rotation $\Omega$, which promotes stability and avoids learning a constant shift of pixels in the prediction.
2.4 Conditional generation of Cresyl violet staining
Image-to-image translation refers to the process of generating images using a generator model $G$ conditioned on an input image $x$. Predictions $G(x)$ of the model are compared with actual target images $y$.
Our style transfer algorithm consists of three main elements:
  1. 1.
    Reconstruction Loss $\mathcal{L}_{R}$ : Similar to Isola et al., (2017), we use an $L_{1}$ loss between target $y$ and prediction $G(x)$ to encourage pixel correspondence as
    $\mathcal{L}_{R}=\mathbb{E}_{x,y}\left\lVert R(y)-G(x)\right\rVert_{1},$ (6)
    where $R$ performs the proposed online registration to spatially align $y$ and $G(x)$ before loss calculation. This enables the utilization of imperfectly aligned training data to penalize any discrepancies in corresponding pixel values between $G(x)$ and $y$.
  2. 2.
    Style Loss $\mathcal{L}_{S}$: We apply a texture-sensitive style loss proposed by Gatys et al., (2015) based on squared distances between Gram matrix representations of neural network features. Given a pre-trained VGG encoder Simonyan and Zisserman, (2015), Gram matrix representations are computed from feature activations of its layers to characterize the texture of images at different complexities. For each layer $l$, the encoder produces a different number of $N_{l}$ feature maps, each storing $K_{l}$ spatial entries (i.e., height $\times$ width). Elements of the Gram matrix $G^{l}_{ij}$ at layer $l$ are computed as the inner product between the $i$-th and $j$-th feature map $F_{i}^{l}$ and $F_{j}^{l}$, where each map is flattened to a $K_{l}$-dimensional vector:
    $G^{l}_{ij}=\sum^{K_{l}}_{k=1}F^{l}_{ik}F^{l}_{jk}.$ (7)
    Since the Gram matrix computation captures global feature correlations rather than spatial locations, this allows a translation-invariant comparison of image statistics. The style loss is computed over all $L$ layers of the VGG encoder, using Gram matrix representations $G^{l}$ for the online registered target and $\hat{G}^{l}$ for the prediction:
    $\mathcal{L}_{S}=\mathbb{E}_{x,y}\left[\sum^{L}_{l=1}\frac{1}{K_{l}^{2}N_{l}^{2}}\sum^{N_{l}}_{i=1}\sum^{N_{l}}_{j=1}(G^{l}_{ij}-\hat{G}^{l}_{ij})^{2}\right].$ (8)
    We refer to this style loss as Gram loss.
  3. 3.
    Equivariance Loss $\mathcal{L}_{E}$: By registering target $y$ to generator prediction $G(x)$ before loss calculation, displacing objects in $G(x)$ relative to its input $x$ is not penalized by reconstruction loss $\mathcal{L}_{R}$. To prevent pixel shifts in $G(x)$, we introduce a third loss component enforcing equivariance with respect to rotations as
    $\mathcal{L}_{E}=\mathbb{E}_{x}\left\lVert\Omega(G(x))-G(\Omega(x))\right\rVert_{2},$ (9)
    where operator $\Omega$ represents an image rotation of 180°180°180  °. Minimizing Eq. 9 ensures that pixels in $G(x)$ correspond to pixels at the same pixel coordinates in $x$ as any discrepancy would cause a mismatch of $\Omega(G(x))$ and $G(\Omega(x))$.
All components are aggregated into total loss
$\mathcal{L}=\lambda\mathcal{L}_{R}+(1-\lambda)\mathcal{L}_{S}+\eta\mathcal{L}_{E},$ (10)
with relative weightings $\lambda\in[0,1]$ and $\eta\geq 0$ as hyperparameters. An overview of the framework is illustrated in Fig. 3.
As an alternative for style loss $\mathcal{L}_{S}$, we consider adversarial training Goodfellow et al., (2014) using a GAN loss to compare with previous work in virtual staining Rivenson et al., (2019); Cross-Zamirski et al., (2022). We implement the GAN loss in the form of a Wasserstein GAN Arjovsky et al., (2017). In contrast to conditional GAN training Mirza and Osindero, (2014); Isola et al., (2017), we do not condition the discriminator on input images $x$, as this would cause the model to reproduce any misalignment in the training data. We refer to models trained with Gram loss for $\mathcal{L}_{S}$ as Gram and models trained with GAN loss as GAN. In combination with the online registration head, we refer to the models as Gram+Reg and GAN+Reg.
2.5 Model training
For the generator $G$, we use the same 5-layer U-Net Ronneberger et al., (2015) with numbers of features [32, 64, 128, 256, 512] in all experiments, and adjust the input and output channels to 3 according to our setup. To train $G$, we use square 3D-PLI patches of 444 pixels size, represented by parameter maps transmittance ($I_{T}$), retardation ($\sin\delta$) and direction ($\varphi$). We reformulate the 3D-PLI parameters as triplets $(I_{T},\sin\delta\cos(2\varphi),\sin\delta\sin(2\varphi))$ to resolve the circular behavior of direction $\varphi$, standardize the channels and stack them to the input of generator $G$. Due to the fully convolutional approach of the U-Net model without padding, the generated output predictions have a reduced size of 260 pixels. Unless specified otherwise, we use a patch size of 360 pixels for the target Cresyl violet images, centered at the input patch position. They are chosen to be larger than the model predictions to allow the online registration to correct translational shifts of up to 50 pixels in any direction, while keeping the predictions fully contained within the target images. We normalize image pixel values of the Cresyl violet staining to the range of [0, 1].
For computing style loss $\mathcal{L}_{S}$, we extract features from a VGG19 model Simonyan and Zisserman, (2015) to compute the Gram loss. The VGG feature encoder network has a depth of four layers and three input channels with pre-trained weights on ImageNet Deng et al., (2009). We multiply the style loss $\mathcal{L}_{S}$ by a constant factor of 104 to bring it to the same order of magnitude as reconstruction loss $\mathcal{L}_{R}$. For the training, we use Adam optimizer Kingma and Ba, (2017) with $\beta_{1}$ = 0.9, $\beta_{2}$ = 0.999 and a learning rate of 10-3. If not stated otherwise, we use $\eta$ = 0.1 and $\lambda$ = 0.5 as default in the training objective in Eq. 10.
In the case of GAN style loss $\mathcal{L}_{S}$, we use a 4-layer convolutional network as discriminator with kernel size 4, stride 2, padding 1, and feature size of [32, 64, 128, 256], followed by batch normalization Ioffe and Szegedy, (2015) and Leaky ReLU after each convolution. For training, we use the Wasserstein GAN Arjovsky et al., (2017) objective and a separate Adam optimizer for the discriminator and the generator, using $\beta_{1}$ = 0.5, $\beta_{2}$ = 0.999 and a learning rate of 10-4. We perform five updates for the discriminator for one update of the generator and clamp discriminator weights at 0.03 after each step.
Figure 4: Localization of train and test data. (A) Seven sections used for training (blue stripes) and one section used for testing (red stripe) were taken at the level of the central sulcus (CS; yellow dashed lines). Locations are shown on top of the 3D reconstructed blockface of the brain for reference. Train and test data are 0.6 mm apart from each other. (B) Selected locations of train and test regions of interest (ROIs), which are used for training and testing the models. The images show ROIs from each of the train and test sections on top of globally affine registered Cresyl violet images. Black contour plots outline the pial surface of corresponding 3D-PLI sections for reference.
Training data
For model training and evaluation, we use eight coronal sections at the level of the central sulcus. Seven sections are used for training and one section is kept for testing with a gap of 0.6 mm between train and test sections (Fig. 4 A). From the training sections, 27 affine-aligned ROIs are extracted (Fig. 4 B), where one ROI is held out for validation to identify possible overfitting. For each ROI, we extract joint target Cresyl violet images and 3D-PLI modalities at the same center location. To maximize the diversity of the training examples, we do not pre-compute training patches but sample them randomly during the training process. We use a batch size of 128 and draw 32,768 paired random patches per epoch evenly distributed across the training ROIs, resulting in approximately 1,260 random samples per ROI per epoch. All training is performed for 150 epochs or until model convergence if validation loss did not decrease for at least 50 epochs.
Data augmentation
To enhance the robustness of our trained models we employ 3D-PLI-specific data augmentations, which were carefully modified to keep physically plausible signal parameters Oberstrass et al., (2024). Specifically, we perform random rotation by angles between -180° and 180° with mirror padding and horizontal and vertical random flipping. In both cases, direction parameter maps $\varphi$ are corrected accordingly. Additionally, we perform Gaussian blurring of 3D-PLI parameter maps for random standard deviations up to $\sigma$ =1.5 and kernel sizes of 3 or 5. We scale thickness and attenuation coefficients for 3D-PLI parameter maps by random values between 0.5 and 2.
Implementation
All models were trained on the supercomputer JURECA-DC at the Jülich Supercomputing Centre (JSC, Thörnig, 2021) on a single node by splitting each batch equally onto 4 NVIDIA A100 GPUs using distributed data-parallel strategy. For data pre-processing 128 worker processes were spawned on 128 CPU cores. For reference, training for 100 epochs took 8 hours with online registration and 4 hours without on this hardware. The implementations are based on the Quicksetup-ai template by the HelmholtzAI Consultants Munich [^1] https://github.com/HelmholtzAI-Consultants-Munich/Quicksetup-ai, using the frameworks: PyTorch Paszke et al., (2019), PyTorch Lightning Borovec et al., (2022) and Hydra Yadan, (2019).
Online registration
For the online registration head, we restrict accepted solutions of Eq. 5 to translation + rotation pairs that cause the registered target to have full overlap with the prediction, avoiding loss calculation over zero-padded values. We check the translation correction for 31 rotation angles from -7.5° to 7.5° with steps of 0.5° and take the translation + rotation pair with the best registration score.
3 Experiments and Results
We compare a selection of performance scores to identify optimal hyperparameters of the proposed method and assess the overall potential of the best performing model. The main hyperparameters are the choice of the online registration metric, the type of style loss $\mathcal{L}_{S}$, its relative weighting $\lambda$ to reconstruction loss $\mathcal{L}_{R}$, and whether to use the additional equivariance loss $\mathcal{L}_{E}$. To reduce the massive computational demands by a rigorous grid search across all hyperparameters, we choose to identify a suitable choice for the online registration metric and model variants using different loss components independently before determining an optimal relative weighting $\lambda$.
3.1 Experiment setup
3.1.1 Test data
We manually select four ROIs for model evaluation, ensuring a diverse representation of different cytoarchitectonic characteristics from the held-out test section (Fig. 4B). The ROIs contain primary motor cortex 4a, temporal cortical area TE, the hippocampal cornu Ammonis (CA) region, and parts of the putamen and globus pallidus (Pars interna and Pars externa) as subcortical structures (Fig. 5).
As the computation of image metrics requires a precise alignment of test data, we perform elastic registration of test ROIs based on landmarks and image intensity using the bUnwarpJ Arganda-Carreras et al., (2006) algorithm. We use predictions of an independently trained Gram+Reg model as target. The predictions are used to manually identify 15-25 characteristic cell clusters as landmarks per ROI, which are confirmed by the location and arrangement of faint shadows of cells visible in the 3D-PLI transmittance. For registration, we use an image weight of 1.0, a landmark weight of 10.0, and a consistency weight of 10.0. The strong weights for landmarks and consistency are chosen to prevent the transformation field from overly conforming to the predictions and to overcome local optima.
Figure 5: Overview of ROIs used for the evaluation, which represent distinct cellular architectures. They were extracted from section 559, located 0.6 mm apart from the training sections. Embedded windows show magnified details inside each ROI. The locations of the magnified views are indicated by yellow frames. Columns each show one of the four test ROIs taken from the anterior subdivision of the primary motor cortex (4a), the hippocampal cornu Ammonis (CA) region, temporal cortical area TE, and parts of the putamen and globus pallidus (GP; a: GP Pars interna; b: GP Pars externa) as subcortical nuclei. The first three rows demonstrate 3D-PLI modalities transmittance $I_{T}$, retardation $\sin\delta$ (scaled using gamma correction for visualization) and fiber orientation in HSV color space (hue: direction $\varphi$; saturation, brightness: retardation $\sin\delta$). The 3D-PLI modalities are compared to the registered target Cresyl violet and predicted virtual staining.
3.1.2 Evaluation scores
We evaluate the impact of different model parameter choices on the quality of the predicted virtual staining by applying structural similarity index measure (SSIM, Wang et al., 2004), mutual information (MI), and root-mean-square error (RMSE). For each metric, we report the mean over all test ROIs. To evaluate the ability to localize cell bodies, we compute F1 scores for instance segmentations of predicted and target cells using a contour proposal network (CPN, Upschulte et al., 2022, 2023). We only consider cells with a minimum in-plane size of 100 µm² and require a minimum intersection over union (IoU) of 30%. Smaller cells could not be reconstructed accurately and would introduce noise into the evaluation. For the cell detection model, we fine-tuned a pre-trained222ginoro_CpnResNeXt101UNet-fbe875f1a3e5ce2c CPN for cell body segmentation in cell-stained microscopy images using the celldetection [^3] https://github.com/FZJ-INM1-BDA/celldetection Python package. Fine-tuning was performed on a diverse mix of manually annotated images, as well as synthetic data.
We restrict the computation of evaluation metrics to gray matter, where most of the neuronal cell bodies are located, excluding white matter and background pixels. In test ROIs showing cortical regions, we further exclude molecular layer I as it consists of only a few individual neurons and exhibits a nonlinear deformation due to tissue shrinkage, which could not be reliably corrected in the registration. For the ROI showing the hippocampus, we exclude the fascia dentata, as its granular layer consists of very densely packed neurons, which cannot reliably be distinguished and restrict the analysis to hippocampal CA1-CA4 regions.
3.2 Online registration metric
Figure 6: Robustness of different online registration metrics against synthetic image distortions. We report the proportion of correct rigid alignments for different blur (A) and noise (B) levels by registration metrics MSE, CC, PC, and BIPC. The registration is considered successful (a hit) if the translational displacement obtained remains within 5 pixels of the actual displacement.
To compute a pixel-aligned reconstruction loss we apply the online registration head (Sec. 2.3). We consider correlation-based registration metrics CC and MSE with phase correlation (PC, Kuglin, 1975) and blur-invariant phase correlation (BIPC, Ojansivu and Heikkila, 2007). We apply a Hann window to PC and BIPC before their calculation to mitigate their bias towards the sharp image edges Gonzales and Woods, (2008).
To understand how image degradation effects can influence the success of the online registration head, we compare the robustness of the metrics against noise and blur. A square target image with 460 pixels size is extracted from a random location in the Cresyl violet staining. Next, a smaller moving tile with a size of 260 pixels is extracted from a random location within the target and distorted by gradually increasing noise or blur on the image. We add noise from a zero-centered Gaussian distribution with standard deviation $\sigma$ increasing from 0 to 25 and Gaussian blur with increasing kernel sizes with standard deviations $\sigma$ growing from 0 to 50. The registration head is applied with each metric to realign the degraded moving tile with its location in the target image. We compare the hit rate over 100 examples per degradation step. The registration is considered successful if the determined translational displacement remains within 5 pixels of the actual displacement.
Fig. 6 shows that as blur increases, BIPC with a Hann window achieves the best performance. This result is expected, as the metric remains invariant to blur. We also observe that CC and MSE perform best in mitigating the effects of noise, while the other metrics fail already with a minor amount of noise. Overall, MSE provides a balanced tradeoff between blur and noise response.
Table 1: Performance of models trained with different online registration metrics in terms of similarity of generated images with ground truth. We compare cross-correlation (CC), phase-correlation (PC), blur-invariant phase-correlation(BIPC), and mean-squared error (MSE) as online registration metrics against using no online registration (-).
Method Metric MI $\uparrow$ RMSE $\downarrow$ SSIM $\uparrow$ F1 $\uparrow$
Gram - 0.126 33.6 0.354 48.8
Gram+Reg CC 0.136 33.0 0.379 55.2
BIPC 0.149 31.5 0.371 54.0
PC 0.185 31.9 0.415 59.4
MSE 0.226 29.8 0.444 63.7
To make an optimal choice of a registration metric for online registration, we train Gram+Reg models with CC, BIPC, PC, and MSE compared to a Gram model without online registration. We use a Cresyl violet target patch size of 360 pixels and predictions of 260 pixels, centered within the target patch. This allows the online registration to correct for a maximum of 50 pixels translation in each dimension. For Gram+Reg models trained with CC, BIPC and the Gram model, $\lambda=0.1$ was chosen as it improved results compared to $\lambda=0.5$ for PC and MSE.
Tab. 1 shows a quantitative comparison of the models. Independent of the metric, the best results are observed when using Gram+Reg with online registration Among all metrics considered, the model with MSE performs best.
3.3 Performance analysis of different model variants
Figure 7: Including equivariance loss $\mathcal{L}_{E}$ in the training improves cell instance overlap of Gram+Reg and GAN+Reg models. (A) 3D-PLI input as transmittance $I_{T}$, retardation $\sin\delta$ (scaled using gamma correction) and fiber orientation map in HSV color space (hue: direction $\varphi$; saturation, brightness: retardation $\sin\delta$) for an example patch from the pyramidal layer of the hippocampal CA1 region. (B) Registered target Cresyl violet staining, a cell segmentation by a CPN model, and the identified target cell instances. (C) Predicted virtual staining of models trained with equivariance loss $\mathcal{L}_{E}$ enabled (✓) or disabled (✗). Cell instances are overlaid with their target cells. Models trained with equivariance loss show less displacement of cells and increased F1 scores for this specific image patch. Since the displacement occurs randomly, a patch and random seeds for the models are selected that show a particularly strong effect.
To justify including equivariance loss $\mathcal{L}_{E}$ (Eq. 9) in the total loss, Fig. 7 illustrates the effect of training Gram+Reg models with this loss component enabled and disabled. With $\mathcal{L}_{E}$ disabled ($\eta$ = 0), cell instances in the model predictions exhibit a relative displacement. Enabling $\mathcal{L}_{E}$ ($\eta$ = 0.1) improves the overlap of cell instances with the original Cresyl violet staining.
Table 2: Effect of the proposed online registration on models trained with different choices of style loss $\mathcal{L}_{S}$. Each variant is trained with online registration head (Reg.) and equivariance loss ($\mathcal{L}_{E}$) enabled or disabled. MSE metric is used for online registration and $\eta$ = 0.1 for weighting of $\mathcal{L}_{E}$. The deviation is reported as standard error over four independent trainings with different random seeds.
$\mathcal{L}_{S}$ Reg. $\mathcal{L}_{E}$ MI $\uparrow$ RMSE $\downarrow$ SSIM $\uparrow$ F1 $\uparrow$
Gram 0.224 $\pm$ 0.003 29.6 $\pm$ 0.6 0.445 $\pm$ 0.003 63.1 $\pm$ 0.5
0.211 $\pm$ 0.004 30.2 $\pm$ 0.4 0.432 $\pm$ 0.004 61.6 $\pm$ 0.4
0.111 $\pm$ 0.012 33.6 $\pm$ 0.8 0.314 $\pm$ 0.023 38.9 $\pm$ 7.6
0.107 $\pm$ 0.007 34.2 $\pm$ 0.3 0.304 $\pm$ 0.015 40.6 $\pm$ 5.7
GAN 0.160 $\pm$ 0.008 30.5 $\pm$ 0.5 0.389 $\pm$ 0.012 50.3 $\pm$ 2.1
0.126 $\pm$ 0.008 33.0 $\pm$ 0.4 0.347 $\pm$ 0.010 43.0 $\pm$ 1.4
0.085 $\pm$ 0.008 38.1 $\pm$ 0.5 0.249 $\pm$ 0.001 21.1 $\pm$ 0.3
0.096 $\pm$ 0.008 38.8 $\pm$ 1.7 0.253 $\pm$ 0.005 22.2 $\pm$ 1.3
To quantify the impact of the loss component $\mathcal{L}_{E}$ on the predictions, we compare the models with $\mathcal{L}_{E}$ enabled and disabled in Tab. 2. We consider Gram and Gram+Reg, as well as GAN and GAN+Reg as model variants. All models use $\lambda=0.5$ expect for the Gram model without online registration, where $\lambda=0.1$ showed better performance. Since displacements of cell instances in the model predictions occur randomly with varying strength, we report mean and standard errors across four independent trainings. Models trained with $\mathcal{L}_{E}$ enabled show the highest correspondence with the target staining. Without online registration, the advantage of including $\mathcal{L}_{E}$ becomes negligible. Furthermore, results in Tab. 2 demonstrate that Gram models outperform their corresponding GAN models in all evaluation metrics. Regardless of the variant used for style loss $\mathcal{L}_{S}$, all models benefit from the online registration.
Figure 8: Comparison of virtual stainings for a patch of temporal area TE. (A) 3D-PLI input visualized as fiber orientation in HSV color space (hue: direction $\varphi$; saturation, brightness: retardation $\sin\delta$) and transmittance $I_{T}$. (B) Registered target Cresyl violet staining for reference with annotations of cortical layers I-VI and white matter (wm). (C) Predicted virtual stainings of GAN and Gram models and extended variants using MSE online registration (GAN+Reg, Gram+Reg). (D) GLI images ranging from the pial surface to the cortex/white matter transition and average profiles (red lines) for layers II - VI.
A comparison of model predictions is shown in Fig. 8C for the whole cortical depth of an Isocortex sample from temporal cortical area TE. GAN+Reg and Gram+Reg provide the most realistic-looking reconstructions in terms of relative size and shape of cell bodies, along with the differences between layers concerning cell packing density. While Gram+Reg seems to be influenced by technically-related inhomogeneities in staining intensity, GAN+Reg produces a clearer contrast between stained cells and surrounding tissue. Gram+Reg detects the transition from the cortex to white matter better than GAN+Reg. Both methods introduce an artificial arrangement of cells into cortical columns not present in the original Cresyl violet staining.
To further analyze the ability of the methods to reconstruct the laminar cell organization, we compute grey level index (GLI) values Schleicher et al., (1978); Zilles et al., (1978). GLI values provide an established proxy for volume density of stained cell bodies in gray matter regions and are used to characterize the laminar architecture of cortical areas Schleicher et al., (2000). To compute GLI values, adaptive thresholding is applied to create a mask of pixels occupied by cell bodies for each image. The masks are subsequently down scaled by a factor of 16 to a spatial resolution of 20.8 µm, with each value representing the fraction of segmented pixels. Resulting GLI images, cropped to the area between pial surface and the cortex/white matter transition, are displayed for the Cresyl violet target and the predictions in Fig. 8D.
For each of the GLI images, 31 intensity profiles vertically oriented to the cortical layers are extracted to represent the columnar distribution of GLI values. The profiles are restricted to layers II - VI since staining inhomogeneities make GLI values for layer I unreliable. To obtain one representative profile for each image, the profiles are averaged and smoothed using a mean filter with kernel size 3 to improve the signal-to-noise ratio. The average profiles are shown on top of each GLI image in Fig. 8D.
The average GLI profile for the GAN model shows no clear distinction between cortical layers II-IV and misses a peak for layer IV. The profile by Gram replicates peaks for layers II and IV but does not provide a clear representation of the cortical layers. GAN+Reg and Gram+Reg both replicate peaks for the higher cell densities in layers II, IV and VI. Gram+Reg shows the clearest distinction between layers and even replicates slightly nuanced peaks of the profile within layers.
3.4 Optimal balancing of style and content
Table 3: Different values for weighting parameter $\lambda$ show that a balance between style loss $\mathcal{L}_{S}$ and reconstruction loss $\mathcal{L}_{R}$ is required in the training of a Gram+Reg model. Setting $\lambda=0$ means only $\mathcal{L}_{S}$ is used. Setting $\lambda=1$ means only $\mathcal{L}_{R}$ is used.
Method $\lambda$ MI $\uparrow$ RMSE $\downarrow$ SSIM $\uparrow$ F1 $\uparrow$
Gram+Reg 0 0.124 33.8 0.340 48.7
0.03 0.162 31.9 0.394 57.7
0.25 0.210 30.7 0.436 62.8
0.5 0.226 29.8 0.444 63.7
0.75 0.238 29.8 0.448 63.2
0.97 0.230 30.7 0.446 59.6
1.0 0.090 38.0 0.347 0.0
Figure 9: Predicted virtual Cresyl violet stainings for different weightings between reconstruction and style loss components. (A) 3D-PLI inputs as transmittance $I_{T}$, retardation $\sin\delta$ and fiber orientation in HSV color space (hue: direction $\varphi$; saturation, brightness: retardation $\sin\delta$). The retardation has been scaled using gamma correction for visualization. (B) Registered target Cresyl violet staining. (C) Predicted virtual stainings of models trained with different weightings $\lambda$ between reconstruction loss $\mathcal{L}_{R}$ and style loss $\mathcal{L}_{S}$. Setting $\lambda=0$ focuses on $\mathcal{L}_{S}$ and setting $\lambda=1$ on $\mathcal{L}_{R}$ exclusively. Gram loss is used for $\mathcal{L}_{S}$. Within each row, yellow markers indicate the same structures across all columns. (i) Temporal area TE and underlying white matter with arrowheads highlighting blood vessels within white matter and cortical layer VI. (ii) Dentate gyrus of the hippocampus with a dashed line delineating the proximal end of the granular layer of the fascia dentata (FD). (iii) Pyramidal layer of the cornu Ammonis (CA1) region of the hippocampus highlighted by dashed lines. (iv) Layer V of the primary motor cortex with arrowheads highlighting two Betz cells.
To identify an optimal weighting $\lambda$ of the style loss in Eq. 10, we train Gram+Reg models with different values for $\lambda$ and MSE online registration metric. A quantitative evaluation of the trained models in Tab. 3 shows that larger values for $\lambda$ up to $\lambda$ = 0.75 achieve the best MI, RMSE, and SSIM scores. This indicates that the reconstruction loss should be given a stronger weight but should not be used alone ($\lambda=1$). The highest overlap of predicted cell instances measured by the F1 score is obtained for a balanced weighting with $\lambda$ = 0.5.
A visual comparison of predictions of Gram+Reg models trained with different choices for $\lambda$ is shown in Fig. 9 for selected crops from the test data. By using $\mathcal{L_{R}}$ only (i.e. $\lambda=1$), the model is unable to reconstruct details from the Cresyl violet staining. It is only able to locate strongly pronounced structures such as the granular layer delineated in (ii). Independent of the style loss weighting, the models do not resolve individual cell instances present within the granular layer. Instead, they capture the overall cell density, represented by a continuous dark purple color. By increasing emphasis on $\mathcal{L}_{S}$ (i.e. with decreasing values for $\lambda$), the generator is able to reconstruct details such as neuronal cell bodies in less dense regions in (i)-(iv) and blood vessels in (i). It is also able to generate Glial cells to match the appearance of white matter in (i) and (iii). However, generators trained with an overweighting of style loss also tend to miss some strongly pronounced structures in the predictions, such as the blood vessels shown in (i) or the Betz cells in (iv).
3.5 Reliability of predicted cell instances across cell sizes
Figure 10: Reliability of predicted cell instances increases with their size. We report F1 scores between detected cell instances by a CPN model in the predicted and target stainings. F1 scores are reported for intervals of the segmented in-plane cell body area. Each bar represents the average over four independently trained models. Error bars show standard deviation.
Larger cells are expected to be more pronounced in 3D-PLI parameter maps compared to smaller cells that may be overshadowed by other tissue components such as nerve fibers. To quantify the reliability of predicted cells in the virtual staining in relation to their size, we compute F1 scores for multiple bins of 2D in-plane cell sizes. Obtaining F1 scores for each bin requires computation of true and false positives, as well as false negatives through matching of predicted and target cells with a minimum IoU threshold of 0.3. We modify the computation of true and false positives by matching only predicted cells within that range with all cells in the target image. To count false negatives, target cells within that range are matched with all cells in the prediction. Cell sizes are corrected with the 2D shrinkage factor of 0.97 estimated in Sec. 2.1.2, assuming homogeneous shrinkage of cells and surrounding tissue.
Results in Fig. 10 show that smaller cells < 50 µm have much lower F1 scores, below 20.0, with all methods. With increasing 2D cell size, they can be identified more accurately. The Gram+Reg model shows highest F1 scores across all cell sizes and smallest variation between four independently trained models, measured as standard deviation.
4 Discussion
We introduced an online registration head capable of approximating smooth nonlinear deformations, given sufficiently small patch sizes for training. Its accuracy is inherently influenced by the choice of the registration metric. We identified MSE in the Fourier domain through multiple applications of cross-correlation as superior to other typically used metrics in Fourier-based image registration, including PC. This is a notable observation, as PC is a popular choice due to its robustness to intensity variations and frequency-dependent noise Tong et al., (2019). However, during early training stages, model predictions are often blurred. The network tends to learn coarse, low-frequency structures first. Additionally, random weight initialization introduces noise in the first steps. MSE appears more resilient to such blur and frequency-independent noise, leading to more accurate predictions (Fig. 6). This behavior can be explained by PC’s reliance on strong high-frequency spectral peaks, which are diminished by blur Pedone et al., (2013). MSE, in contrast, maintains sensitivity across the frequency spectrum and can better handle smooth, low-detail inputs that dominate during initial training.
The applied pixel-wise reconstruction loss assumes a precise alignment to provide an informative feedback signal. However, the online registration head will typically only succeed if a good contrast between structures is being predicted. In other words, there is a cross-dependency of accurate pixel reconstructions and producing a high contrast over training iterations, which may result in a conflict during training. This can be seen in Fig. 9, where a Gram+Reg model focussing on pixel reconstruction only made heavily blurred predictions that do not provide structural details for cell-precise registration. Adding a style loss component leads to more pronounced contours while simultaneously achieving higher pixel reconstruction accuracy (Tab. 3) through online registration. This underpins the importance of the additional style loss, which does not depend on perfect alignment and guides the training to produce structures.
Apparently, the choice for Gram or GAN as style loss should be considered in the light of the application. Training with Gram loss led to more accurate cell localization and pixel values (Tab. 2). However, it was affected by staining intensity inhomogeneities, resulting in blurry cells (Fig. 8). GAN loss, on the other hand, produced higher contrast in cells but with lower accuracy. This could be a consequence of the Gram loss referring to the specific target image, while the GAN loss refers to the average scoring of the discriminator across the distribution, less than the precision with respect to an individual image.
The choice of style loss function also impacts the stability and complexity of the training process. When using GAN loss, the training process has to be carefully monitored. If training collapses, restarting is often necessary. It requires careful balancing of discriminator and generator capacities and tuning of style and reconstruction loss weighting. Training with Gram loss, on the other hand, is generally more convenient to train. It requires no balancing of model capacities, is robust to loss balancing, and does not collapse during training.
Throughout all models, we observed occasional staining inhomogeneities in the predicted virtual stainings. This effect was particularly pronounced in the form of stripe patterns in the primary motor cortex ROI (Fig. 5). Cell body membranes and blood vessel walls pose hard edges in the cortex, resulting in diffraction patterns even brighter than the background field. Varying of the focus level of the LMP-1 microscope allows optical scanning of the Poisson spots induced by diffraction. Cell bodies are highlighted with higher contrast by shifting and lowering the focal plane to the top and the bottom of the tissue by approximately $\pm$30 µm (Fig. 11). This means that the ability to predict cells from 3D-PLI measurements is sensitive to the focus level of the microscope, and it is difficult to generalize to unseen focus levels that were not part of the model training.
During the mounting process of the 60 µm thick sections in 20% glycerol solution on the object slide the gray matter is more prone to swelling than the more compact white matter, leading to tissue expansion, mechanical stress and tape flutter especially in the cortex due to its higher flexibility. The free-floating tissue between object slide and coverslip shapes waves of up to 30 µm in amplitude which affects the focus level and the light incidence on the tissue, hence the effective birefringence, modifying the polarization state of the transmitted light. Therefore, the stripe-like inhomogeneities in the virtual staining can be explained by such variations in focus levels within the same section.
An effective countermeasure against waves in the tissue is to weight the open coverslips after embedding overnight. In addition, the acquisition of multiple focus levels for each section could potentially provide missing information. For the given dataset, however, these potential solutions were not applicable, and measurements could not be easily repeated. In addition, the acquisition of multiple focus levels would significantly increase the measurement effort. Nevertheless, it will be an important consideration for future investigations to include multiple focus levels in each measurement to improve cell extraction.
While most of the larger cells (>100 µm² in-plane size) could be localized in 3D-PLI parameter maps, smaller cells were often missed or misplaced in the predictions. This is expected as smaller cell structures are often dominated by a stronger signal of intersecting nerve fibers or overlay each other due to the high section thickness of 60 µm required for 3D-PLI. The predictions also showed a stronger arrangement of cells into cortical columns than in the target images. The presence or absence of this arrangement is one of the criteria used to identify cortical areas, which might impair downstream interpretation of the predicted architecture. It may be attributed to the pronounced signal of radial fibers in 3D-PLI, which are used by the virtual staining model to infer potential cell locations. Further investigations are necessary to understand this effect.
Our ablation experiments in Sec. 3.2, Sec. 3.3 and Sec. 3.4 support the design decisions of the proposed virtual staining method. They showed an overall good reconstruction and localization of individual cell bodies. But, of course, the virtual stainings can not yet fully replace real measurements. They occasionally include staining inhomogeneities, miss smaller cells, or introduce implausible cytoarchitectonic features, such as artificial columnar arrangements of cells. In contrast, extracting GLI profiles from the predictions demonstrated that the laminar cell organization can be reconstructed quite well. However, further investigations with larger datasets spanning multiple brain areas are necessary to thoroughly assess their reliability.
The proposed methodology might be further improved by extending the amount of training sections, brain regions and focus levels in 3D-PLI acquisition. We also expect that the expressiveness of the Gram matrix representation could be enhanced by replacing the VGG encoder, pre-trained on ImageNet, with a domain-specific model, which is powerful on microscopy data. This is technically motivated by the observation that higher-level features in ImageNet-pretrained networks, such as VGG, are optimized for object semantics in natural images (e.g., faces, animals, vehicles), which are irrelevant in the context of fine-grained texture characteristic of microscopy images Stuckner et al., (2022). In contrast, microscopy-specific encoders are better suited to capture such low- to mid-level textural features and have been shown to improve clustering Oberstrass et al., (2024), segmentation Stuckner et al., (2022), and the evaluation of generative models Kropp et al., (2024) over those trained on ImageNet. Therefore, computation of Gram matrices using domain-specific encoders Spitzer et al., (2017); Schiffer et al., 2021b ; Schiffer et al., 2021a or emerging foundation models for histology Tran et al., (2024) might further enhance representation of cytoarchitectural characteristics.
5 Conclusions
Motivated by previous observations that larger cells are encoded in 3D-PLI parameter maps alongside fiber orientations Zeineh et al., (2017), we introduced a deep learning model for transforming 3D-PLI maps into virtual Cresyl violet cell body stainings. This approach enables joint visualization of fiber tracts and cell bodies in the same tissue. Compared to real post-staining, the model offers a scalable alternative that avoids manual labor, limited only by its generalization to unseen architectural patterns. Validation against post-stained test sections confirms that our virtual stainings reliably localize most of the larger cell bodies (>100 µm² in-plane size).
A central contribution of our approach is the integration of an online registration head during training. This component eliminates the need for explicit multimodal registration, which is commonly required in virtual staining pipelines Rivenson et al., (2020); de Haan et al., (2021); Yang et al., (2022). It is a simple but highly efficient add-on that can be combined with various loss formulations, leveraging model-estimated landmarks, to continuously refine the alignment over time.
The method enables accurate prediction of cell bodies from 3D-PLI and successful adaptation to the appearance of real Cresyl violet stainings. As such, it expands the usability of 3D-PLI in large-scale data settings, allowing virtual staining at scale. While such synthetic data cannot and should not replace real histological measurements, it offers promising opportunities in downstream analysis. Potential applications include cross-modal image registration to align real Cresyl violet and 3D-PLI images, disentangling signal sources to identify voxels dominated by cell bodies (e.g., for improved tractography), and missing data imputation in serial section stacks. Especially in interleaved modalities, this could enable the reconstruction of complete datasets. Of course, such applications require careful quality control and distinction between synthetic and real measurements.
The outcomes of this study lay the groundwork for prospective investigations focused on enhancing 3D-PLI analysis, particularly through the exploration of dedicated cell detection techniques to directly extract cell body instances from 3D-PLI data. Furthermore, the findings serve as a motivation for gathering additional training data, aiming to refine and extend the application of the virtual Cresyl violet staining to a broader range of sections, brains, and species. While the current model is trained specifically to replicate Cresyl violet stainings, the same methodology can in principle be adapted to other staining types, provided appropriate retraining is performed. Future research should focus on further investigating its transferability to other datasets, staining protocols, and brains.
Ethics Statement
Vervet monkeys used were part of the Vervet Research Colony housed at the Wake Forest School of Medicine. Our study did not include experimental procedures with living animals. Brains were obtained when animals were sacrificed to reduce the size of the colony, where they were maintained and sacrificed in accordance with the guidelines of the Wake Forest Institutional Animal Care and Use Committee IACUC #A11-219 and the AVMA Guidelines for the Euthanasia of Animals.
Data and Code Availability
The training pipeline for presented Gram+Reg and GAN+Reg models is available on GitHub [^4] https://github.com/FZJ-INM1-BDA/pli2cells.
Code for the online registration  [^5] https://jugit.fz-juelich.de/inm-1/bda/software/data_processing/ffreg, data augmentations for 3D-PLI images  [^6] https://jugit.fz-juelich.de/inm-1/bda/software/data_processing/pli-transforms, visualization methods for 3D-PLI modalities  [^7] https://jugit.fz-juelich.de/inm-1/bda/software/data_processing/pli-styles, as well as additional dependencies  [^8] https://jugit.fz-juelich.de/inm-1/bda/software are hosted on our external GitLab server.
The ROIs employed for the training and testing of the models in this study, along with a selection of model predictions, are available in our central institutional repository Oberstrass et al., (2025).
Author Contributions
Alexander Oberstrass: Conceptualization, Methodology, Software, Formal analysis, Investigation, Data curation, Writing - original draft, Visualization. Esteban Vaca: Conceptualization, Methodology, Software, Formal analysis, Investigation, Data curation, Writing - original draft, Visualization. Eric Upschulte: Software, Formal analysis, Writing - review & editing, Visualization. Meiqi Niu: Validation, Data curation, Investigation, Writing - review & editing. Nicola Palomero-Gallagher: Validation, Investigation, Writing - review & editing. David Graessel: Validation, Investigation, Writing - review & editing. Christian Schiffer: Conceptualization, Methodology, Software, Writing - review & editing. Markus Axer: Conceptualization, Validation, Investigation, Supervision, Writing - review & editing, Project administration, Funding acquisition. Katrin Amunts: Conceptualization, Validation, Investigation, Supervision, Writing - Review & Editing, Resources, Project administration, Funding acquisition. Timo Dickscheid: Conceptualization, Methodology, Supervision, Writing - original draft, Writing - Review & Editing, Project administration, Funding acquisition.
Conflict of Interest Statement
The authors declare that the research was conducted in the absence of any commercial or financial relationships that could be construed as a potential conflict of interest.
Funding
This project received funding from the Helmholtz Association’s Initiative and Networking Fund through the Helmholtz International BigBrain Analytics and Learning Laboratory (HIBALL) under the Helmholtz International Lab grant agreement InterLabs-0015, the Helmholtz Association portfolio theme “Supercomputing and Modeling for the Human Brain”, the Priority Program 2041 (SPP 2041) "Computational Connectomics" of the German Research Foundation (DFG), the Federal Ministry of Education and Research (BMBF) under project number 01GQ1902, the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation; PA 1815/1-1), and the European Union’s Horizon 2020 Research and Innovation Programme, grant agreement 945539 (HBP SGA3), which is now continued in the European Union’s Horizon Europe Programme, grant agreement 101147319 (EBRAINS 2.0 Project). Computing time was granted through JARA on the supercomputer JURECA at Jülich Supercomputing Centre (JSC). Vervet monkey research was supported by the National Institutes of Health under grant agreements R01MH092311 and P40OD010965.
Acknowledgments
The authors gratefully thank Roger Woods (David Geffen School of Medicine, UCLA, USA) and Karl Zilles (INM-1, Forschungszentrum Jülich GmbH, Germany) for their collaboration in the vervet brain project, and the lab team of the INM-1 for preparing the brain sections. We also sincerely thank the members of the Fiber Architecture and Big Data Analytics groups of the INM-1 for their constructive discussions and thoughtful insights throughout the development of this work.
Appendix A Appendix
Figure 11: Virtual staining quality varies with the focus level of the LMP-1 due to differences in the captured diffraction patterns. Example images show the vervet entorhinal cortex with zoom-ins to the architecture of cortical layers II-III. 3D-PLI inputs are represented by the transmittance $I_{T}$, retardation $\sin\delta$ (scaled with gamma correction) and fiber orientation in HSV color space (hue: direction $\varphi$; saturation, brightness: retardation $\sin\delta$). The same Gram+Reg model is used to predict virtual stainings from a focus level manually adjusted to the tissue center (+0 µm) and 30 µm above and below. A lower focus level of -30 µm achieves the highest consistency in staining intensity and produces the most cells. Higher focus levels introduce an increasing amount of staining artifacts.

### v2

@@
 3D-PLI acquisition is label-free and allows subsequent staining of sections after 3D-PLI measurement.
-By post-staining for cell bodies, a direct link between fiber- and cytoarchitecture can potentially be established within one and the same section.
+By post-staining for cell bodies, a direct link between fiber- and cytoarchitecture can potentially be established in the same section.
 However, inevitable distortions introduced during the staining process make a costly nonlinear and cross-modal registration necessary in order to study the detailed relationships between cells and fibers in the images.
-In addition, the complexity of processing histological sections for post-staining only allows for a limited number of samples.
+In addition, the complexity of processing histological sections for post-staining only allows for a limited number of such samples.
 In this work, we take advantage of deep learning methods for image-to-image translation to generate a virtual staining of 3D-PLI that is spatially aligned at the cellular level.
 We use a supervised setting, building on a unique dataset of brain sections, to which Cresyl violet staining has been applied after 3D-PLI measurement.
-To ensure a high correspondence between both modalities, we address the misalignment of training data using Fourier-based registration methods.
+To ensure high correspondence between both modalities, we address the misalignment of training data using Fourier-based registration.
 In this way, registration can be efficiently calculated during training for local image patches of target and predicted staining.
-We demonstrate that the proposed method enables prediction of a Cresyl violet staining from 3D-PLI, matching individual cell instances.
+We demonstrate that the proposed method can predict a Cresyl violet staining from 3D-PLI, resulting in a virtual staining that exhibits plausible patterns of cell organization in gray matter, with larger cell bodies being localized at their expected positions.
 Keywords deep learning, virtual staining, fiber architecture, cytoarchitecture, polarized light imaging, Cresyl violet, vervet monkey brain
@@
 To understand the organizational principles of the brain, complementary imaging techniques are used to highlight different aspects of brain architecture.
-Two important aspects of the microstructural organization are fiber- and cytoarchitecture Nieuwenhuys, (2013); Amunts and Zilles, (2015).
+Two important aspects of the microstructural organization are fiber- and cytoarchitecture nieuwenhuys2013; amunts2015.
 While cytoarchitecture encompasses the spatial distribution and shape of cell bodies in the cerebral cortex and subcortical nuclei, fiber architecture refers to the course and composition of nerve fibers.
 However, cytoarchitecture and fiber-architecture are usually studied using different staining protocols, applied in different sections.
-Only a few protocols are available to combine cyto- and fiber staining in a single protocol, e.g., Luxol fast blue Klüver and Barrera, (1953), Bielschowsky Bielschowsky, (1904) or the triple staining by Novotny Novotny and and Novotny, (1977) and Nowotny (2009).
-While they allow visualizing cell bodies and fibers in one and the same section, they do not allow the tracing of axons and fiber bundles over long distances.
+Only a few protocols are available to combine cyto- and fiber staining in a single protocol, e.g., Luxol fast blue kluver1953, Bielschowsky bielschowsky1904 or the triple staining by Novotny novotny1977.
+While they allow visualizing cell bodies and fibers in one and the same section, they lack information about 3D fiber orientations.
+As a result, they do not support the tracing of axons and fiber bundles over long distances.
 3D-Polarized Light Imaging (3D-PLI) addresses this limitation.
-It is a microscopic imaging technique for evaluating the three-dimensional orientation of myelinated nerve fibers in entire, unstained histological brain sections Axer and Amunts, (2022); Axer et al., 2011a ; Axer et al., 2011b .
+It is a microscopic imaging technique for evaluating the three-dimensional orientation of myelinated nerve fibers in entire, unstained histological brain sections axer2022; axer2011; axer2011a.
 The technique can achieve an in-plane resolution of 1.3 µm, capturing structures at the level of individual fibers and small fiber bundles.
-3D-PLI has been used to gain insights into the architecture of nerve fibers in different brain regions, such as the human hippocampus Zeineh et al., (2017), the sagittal stratum Caspers et al., (2022) and the vervet monkey visual system Takemura et al., (2020).
-In addition, 3D-PLI has been used to validate fiber tractography algorithms and the interpretation of DW-MRI Caspers and Axer, (2019).
-3D-PLI potentially allows joint imaging of fiber tracts and neuronal cell bodies Zeineh et al., (2017) due to scattering effects affecting light transmittance Menzel et al., (2020).
+3D-PLI has been used to gain insights into the architecture of nerve fibers in different brain regions, such as the human hippocampus zeineh2017, the sagittal stratum caspers2022 and the vervet monkey visual system takemura2020.
+In addition, 3D-PLI has been used to validate fiber tractography algorithms and the interpretation of DW-MRI caspers2019.
+3D-PLI potentially allows joint imaging of fiber tracts and neuronal cell bodies zeineh2017 due to diffraction patterns, differences in the density of birefringent material, and locally variable attenuation.
 However, this possibility has not yet been validated.
@@
 This allows to study cell shape, density and distribution, which vary between brain regions.
-Due to its high spatial resolution, microscopic analysis of histological sections is considered the gold standard to verify structural parcellations Amunts and Zilles, (2015).
-Recent advances in high-throughput scanning, data processing algorithms, and computational capacities have enabled the creation of 3D human brain atlases based on cytoarchitecture, such as BigBrain Amunts et al., (2013), the Allen Adult Human Brain Atlas (AAHA, Ding et al., 2016), and the Julich Brain probabilistic atlas Amunts et al., (2020).
+Due to its high spatial resolution, microscopic analysis of histological sections is considered the gold standard to verify structural parcellations amunts2015.
+Recent advances in high-throughput scanning, data processing algorithms, and computational capacities have enabled the creation of 3D human brain atlases based on cytoarchitecture, such as BigBrain amunts2013, the Allen Adult Human Brain Atlas (AAHA, ding2016), and the Julich Brain probabilistic atlas amunts2020.
 Since 3D-PLI relies solely on optical properties of the tissue, it is label-free and can be combined with a staining of the same tissue after its measurement.
@@
 However, 3D-PLI and Cresyl violet stained tissue share only a few automatically identifiable cross-modal registration landmarks, such as blood vessels or morphological landmarks.
-I.e., post-staining of sections imaged with 3D-PLI is feasible, but technically challenging, and it hardly scales up.
+Therefore, post-staining of sections imaged with 3D-PLI is feasible but technically challenging.
+Since it does not scale efficiently to larger datasets, it is not applicable to whole-brain stacks involving thousands of sections.
 Figure 1: The proposed virtual staining workflow.
@@
 Therefore, we aim to train a deep neural network model to perform image-to-image translation from 3D-PLI to a Cresyl violet staining.
-Such an approach is often denoted as virtual histological staining, which refers to computational methods that generate color-coded images of biological tissue without the need for traditional staining techniques.
+Such an approach is often denoted as virtual histological staining, which refers to computational methods that generate color-coded images of biological tissue without the need for traditional staining techniques latonen2024.
 The methods instead utilize optical properties of the tissue, such as birefringence, autofluorescence, scattering, or absorption to create images that emulate the appearance of stained tissue.
-A virtual Cresyl violet staining spatially aligned with 3D-PLI would allow an application of established tools for cytoarchitectonic analysis, such as automatic cell instance segmentation Upschulte et al., (2022), for 3D-PLI measurements.
-Furthermore, it would provide detailed registration landmarks for cross-modal registration Ounkomol et al., (2018), thereby offering the opportunity to perform joint acquisition of aligned fiber and cytoarchitecture at a larger scale than possible today.
-One of the earliest applications of label-free imaging for visual staining was Quantitative Phase Imaging (QPI, Curl et al., 2004).
+A virtual Cresyl violet staining spatially aligned with 3D-PLI would allow the use of established tools for cytoarchitectonic analysis, such as automatic cell instance segmentation upschulte2022, directly on 3D-PLI data.
+Furthermore, identification of cell bodies would provide detailed registration landmarks for cross-modal registration ounkomol2018, thereby offering the opportunity to perform joint acquisition of aligned fiber and cytoarchitecture at a larger scale than possible today.
+A virtual staining could, in principle, be applied to the whole brain.
+However, white matter regions remain challenging due to the predominance of glial cells, mainly oligodendrocytes forming the myelin sheath, which are not distinguishable from nerve fibers in the 3D-PLI signal.
+Therefore, in the present study, we focus our analysis on gray matter regions.
+One of the earliest applications of label-free imaging for visual staining was Quantitative Phase Imaging (QPI, curl2004).
 QPI measures the phase shift of light passing through a sample, producing high-resolution images that reveal the optical properties of the tissue.
-It was used to generate images of collagen fibers, red blood cells, and other tissue structures without staining Curl et al., (2004); Park et al., (2018).
+It was used to generate images of collagen fibers, red blood cells, and other tissue structures without staining curl2004; park2018.
 Later, machine-learning algorithms, especially generative models, have been trained to recognize and virtually stain different tissue structures in unstained images by performing image-to-image translation.
-They have successfully generated color-coded images of tissue that replicate the appearance of histological stainings, such as a virtual hematoxylin and eosin (H&E), Masson’s trichrome, and Jones’ stain from QPI of label-free tissue Rivenson et al., (2019), a transformation of H&E stained tissue into Masson’s trichrome, periodic acid-Schiff (PAS), or Jones’ stain de Haan et al., (2021); Yang et al., (2022).
+They have successfully generated color-coded images of tissue that replicate the appearance of histological stainings, such as a virtual hematoxylin and eosin (H&E), Masson’s trichrome, and Jones’ stain from QPI of label-free tissue rivenson2019, a transformation of H&E stained tissue into Masson’s trichrome, periodic acid-Schiff (PAS), or Jones’ stain dehaan2021; yang2022.
 However, these stains are not very good at distinguishing the different components of the nervous tissue.
-Machine learning algorithms were also used to predict fluorescence-labeled images from transmitted-light z-stacks Ounkomol et al., (2018); Christiansen et al., (2018); Cross-Zamirski et al., (2022) or 3D fluorescence structures and a FluoroMyelin stain from bright-field and polarization images of brain slices Guo et al., (2020).
-The methods typically use cross-entropy Christiansen et al., (2018), mean absolute (L1) loss Guo et al., (2020), or style-related losses such as conditional generative adversarial network (GAN) loss Rivenson et al., (2019); de Haan et al., (2021); Cross-Zamirski et al., (2022).
-While including a GAN objective encourages prediction of realistic-looking images, it has no clear mechanism to preserve content when conditioned on a particular input image, and thus may introduce artificial structures Cohen et al., (2018).
+Machine learning algorithms were also used to predict fluorescence-labeled images from transmitted-light z-stacks ounkomol2018; christiansen2018; cross-zamirski2022 or 3D fluorescence structures and a FluoroMyelin stain from bright-field and polarization images of brain slices guo2020.
+The methods typically use cross-entropy christiansen2018, mean absolute (L1) loss guo2020, or style-related losses such as conditional generative adversarial network (GAN) loss rivenson2019; dehaan2021; cross-zamirski2022.
+While including a GAN objective encourages prediction of realistic-looking images, it has no clear mechanism to preserve content when conditioned on a particular input image, and thus may introduce artificial structures cohen2018.
 A combination with a pixel-wise reconstruction loss (e.g.
-L1 loss) mitigates this problem of GAN training and improves accuracy of predictions Isola et al., (2017).
-Since methods using paired training data for supervised image-to-image translation typically produce more accurate predictions than unpaired methods Zhu et al., (2017), a pixel-accurate alignment of training data is necessary.
+L1 loss) mitigates this problem of GAN training and improves accuracy of predictions isola2017.
+Since methods using paired training data for supervised image-to-image translation typically produce more accurate predictions than unpaired methods zhu2017; latonen2024, a pixel-accurate alignment of training data is desired.
 This requires virtual staining methods to either perform a costly registration step or directly acquire paired images.
@@
 Therefore, to alleviate the need for perfectly paired training data, we propose a supervised learning objective performing local online registration of training pairs combined with a translational-invariant style comparison.
-This allows us to train the model on imperfectly registered image pairs with strong content preservation as in paired image-to-image translation, while enabling realistic prediction of subtle structures like cell bodies (Fig.
-1).
-The main contributions are as follows:
-  • We apply the matching of Gram matrix representations as a texture sensitive style loss for the virtual staining, as previously used for texture synthesis Gatys et al., (2015).
+This allows us to train the model on imperfectly registered image pairs with strong content preservation as in paired image-to-image translation, while enabling realistic prediction of subtle structures like cell bodies (Fig.˜1).
+The main contributions of our method are the following:
+  • We apply the matching of Gram matrix representations as a texture sensitive style loss for the virtual staining, as previously used for texture synthesis gatys2015a.
 Since the computation of Gram matrices is translation invariant, it allows a direct comparison of image statistics between coarsely aligned training examples.
 It therefore improves the accuracy of predicted cell instances over commonly used GAN style loss.
-  • An online registration head for improving registration accuracy of local image pairs after pre-alignment of larger tissue sections during training.
+  • An online registration head for improving registration accuracy of local image pairs after pre-alignment of larger tissue tiles during training.
 We consider Fourier-based registration methods, which can be computed efficiently in real-time on modern GPU hardware.
@@
 We demonstrate the proposed virtual staining approach on a set of brain sections for which Cresyl violet staining has been performed after 3D-PLI measurement.
-The brain sample for this study was obtained from a healthy 2.4-year-old adult male vervet monkey (Wake Forest-ID 1818, Axer et al., 2020; Takemura et al., 2020; Menzel et al., 2022) in accordance with the Wake Forest Institutional Animal Care and Use Committee (IACUC #A11-219) and conforming the AVMA Guidelines for the Euthanasia of Animals.
+The brain sample for this study was obtained from a healthy 2.4-year-old adult male vervet monkey (Wake Forest-ID 1818, axer2020vervet; takemura2020) in accordance with the Wake Forest Institutional Animal Care and Use Committee (IACUC #A11-219) and conforming the AVMA Guidelines for the Euthanasia of Animals.
 To obtain an undistorted volumetric reference, a T2 weighted MRI was acquired in-vivo one day prior to sacrifice.
@@
 Blockface images of the frozen tissue block were taken with a CCD camera before cutting each brain section.
-The images were reconstructed into a 3D blockface volume to provide an undistorted reference for section realignment.
+The images were reconstructed into a 3D blockface volume to provide an undistorted reference for section realignment schober2015.
 2.1.1 Image acquisition
 Figure 2: Data modalities and registration challenges for training section 544.
-(A-C) 3D-PLI parameter maps: Transmittance $I_{T}$, retardation $\sin\delta$ and fiber orientation in HSV color space (hue: direction $\varphi$; saturation, brightness: retardation $\sin\delta$).
+(A-C) 3D-PLI parameter maps: Transmittance, retardation and fiber orientation in HSV color space (hue: fiber direction; saturation/brightness: retardation).
 Background pixels are masked for visualization purposes only.
@@
 Yellow arrows indicate blood vessels that can be used as mutual registration landmarks for coarse alignment.
-For 3D-PLI acquisition, brain sections were scanned using a polarizing microscope (LMP-1, Taorad, Germany) with 1.3 µm resolution Axer et al., 2011b ; Axer and Amunts, (2022).
+For 3D-PLI acquisition, brain sections were scanned using a polarizing microscope (LMP-1, Taorad, Germany) with 1.3 µm resolution axer2011a; axer2022.
 The focus level of the LMP-1 was manually adjusted to the center of the tissue for each section.
@@
 $I_{\rho}=\frac{I_{T}}{2}\left(1+\sin(2\rho-2\varphi)\sin\delta\right).$ (1)
-Using harmonic Fourier analysis, parameter maps of transmittance ($I_{T}$), retardation ($\sin\delta$), and direction ($\varphi$) were obtained from the measurements, with a resolution of approximately 34,000 $\times$ 44,000 pixels per section, revealing their fine-grained nerve fiber architecture (Fig.
-2A-C).
+Using harmonic Fourier analysis, parameter maps of transmittance ($I_{T}$), retardation ($\sin\delta$), and fiber direction ($\varphi$) were obtained from the measurements, with an image size of approximately 34,000 $\times$ 44,000 pixels per section, revealing their fine-grained nerve fiber architecture (Fig.˜2A-C).
+Each parameter map was stored in a separate HDF5 file as uncompressed 32-bit floating-point single-channel image.
 After 3D-PLI acquisition, brain sections were washed, fixed and stained for cell bodies with Cresyl violet Nissl staining to reveal their cellular architecture.
-They were scanned using flatbed scanners at 1 µm in-plane resolution and saved as RGB color pictures with eight bit color depth, storing pixel values from 0 to 255 (Fig.
-2D).
-2.1.2 Tissue shrinkage correction
-Deformations induced by histological processing include a volumetric shrinkage of brain tissue Amunts et al., (2005).
-To correct this effect, the extent of shrinkage is usually estimated from the ratio between the histological processed and true brain volume represented by the fresh weight of the whole brain Amunts et al., (2005, 2013) or an MRI reference Wagstyl et al., (2020).
-To estimate the shrinkage of the 3D-PLI acquisition, a linear registration of the 3D reconstructed blockface volume and the postmortem MRI of the same brain was performed.
-The determinant of the transformation matrix showed a global volume change below 1%, which can be neglected.
-In a second step, we calculated 2D shrinkage factors for each brain section considered in this work as the quotient between the area occupied by tissue in the 3D-PLI measurement and its corresponding blockface image.
-For test section #559, we estimated a 2D shrinkage factor of 0.97 meaning a slight expansion of the 3D-PLI measurement relative to its original area within the MRI.
-This observation is consistent throughout all sections with 2D shrinkage factors between 0.95 and 0.99.
-We report in-plane cell sizes measured with 3D-PLI multiplied by an individual 2D shrinkage factor per section, assuming a uniform area change of cells and surrounding tissue.
+Whole-slide flat scans (single-plane) were performed using a Huron TissueScope LE120 high-throughput scanner at 1 µm in-plane resolution (Fig.˜2D).
+The resulting images were saved as RGB color images with eight bit color depth (pixel values ranging from 0 to 255) in uncompressed BigTIFF format.
+2.1.2 Optical effects of cell bodies on the 3D-PLI signal
+While 3D-PLI was primarily developed to map fiber orientations, cell bodies contribute to the measured signal as well.
+In the following, we summarize how absorption, diffraction, birefringence and scattering effects of cells are represented in 3D-PLI parameter maps.
+Previous work reported that larger cell bodies appear as dark spots in transmittance maps zeineh2017, which encode light extinction caused by any material along the optical path.
+However, in the present transmittance maps, cell bodies are not distinguished by higher absorption relative to the surrounding fiber architecture (Fig.˜2A).
+This is likely because their membranes remained intact due to the short postmortem time before tissue fixation.
+Diffraction significantly impacts the 3D-PLI signal at the given wavelength and resolution.
+In cortical regions, diffraction can cause pixels of transmittance maps inside the tissue to appear brighter than the background, particularly along sharp edges such as the walls of cell bodies and blood vessels.
+The intensity of these diffraction patterns on the transmittance map depends on the level of the focal plane of the objective lens.
+Another relevant effect is observed in retardation maps, which encode the average amount and orientation of birefringent material within each tissue voxel, primarily collagen and myelinated nerve fibers.
+Since cell bodies contain significantly less birefringent material than surrounding fibers, their presence causes local attenuation of the retardation appearing as dark patches between cortical fibers (Fig.˜2B).
+Light scattering is ubiquitous in polarized light bright-field transmission microscopy, but can generally be treated as random background noise.
+It does not appear to systematically affect either transmittance or retardation, with the only known exception being increased scattering at steep fibers running approximately perpendicularly to the section plane and darkening the transmittance significantly.
+2.1.3 Tissue shrinkage estimation
+Deformations induced by histological processing include shrinkage or swelling of brain tissue.
+To correct these effects, the extent of shrinkage can be estimated from the ratio between the histologically processed and true brain volume, which can be represented by the fresh weight of the whole brain with an estimated mean specific density amunts2005; amunts2013 or an MRI reference wagstyl2020.
+In this work, we perform 2D segmentation of cell bodies in 3D-PLI parameter maps and measure their in-plane areas.
+To ensure comparability of cell sizes across studies, we estimate 2D shrinkage factors using a postmortem MRI of the same brain as a reference.
+We first estimate shrinkage in the 3D-PLI acquisition by affine registration of the 3D reconstructed blockface volume to the MRI.
+The linear part of the affine transformation in physical space has eigenvalues [0.984, 1.005, 1.015], indicating a global volume change of less than 1% and no axis-specific systematic deviation.
+In a second step, we calculate 2D shrinkage factors for each brain section as the quotient between the area occupied by tissue in the 3D-PLI measurement and its corresponding blockface image.
+For test section #559, we estimate a global 2D shrinkage factor of 0.97, indicating a slight swelling of the 3D-PLI measurement relative to its original area within the MRI.
+This observation is consistent across all sections, with 2D shrinkage factors between 0.95 and 0.99.
+We correct the in-plane sizes of segmented cell body areas in 3D-PLI by applying the individual 2D shrinkage factor of each section, assuming an approximately uniform area change of cells and surrounding tissue.
 2.2 Initial cross-modality alignment
@@
 To align both modalities, an initial affine registration of whole brain sections is performed by manual identification of large blood vessels as landmarks visible in both modalities.
-Performing the initial affine registration reveals remaining nonlinear deformations as shown in Fig.
-2D.
+Performing the initial affine registration reveals remaining nonlinear deformations as shown in Fig.˜2D.
 Since nonlinear deformations typically have low spatial frequencies, causing smooth, large-scale distortions, we expected near-linear deformations at smaller scales.
@@
 We subsequently crop square regions of interest (ROIs) with a size of 4,096 pixels ($\sim$5.3 mm) and without visible artifacts, covering distinct cellular architectures across the whole coronal plane.
-For all ROIs we perform additional affine registration and make sure that transformed landmarks have a maximum distance of 70 pixels (91 µm) from their matches.
+For all ROIs we perform additional affine registration and make sure that transformed landmarks have a maximum distance of 70 pixels (91 µm) from their matches (Fig.˜1A).
 All ROIs are warped and resampled to 1.3 µm using linear interpolation to match the coordinate space of 3D-PLI.
@@
 2.3 Fourier-based online registration of image patches
-To correct the remaining misalignment of 3D-PLI and Cresyl violet after affine registration of ROIs at a finer local scale, we introduce an online registration head that performs cross-modality alignment during training based on model predictions of small image patches.
+To correct the remaining misalignment of 3D-PLI and Cresyl violet after affine registration of ROIs at a finer local scale, we introduce an online registration head that performs cross-modality alignment during training based on model predictions of small image patches (Fig.˜1B).
 We assume that once the style transfer model has learned to reconstruct microscopic landmarks (e.g., individual cells or small blood vessels), such online registration will promote the learning of additional landmarks until the training can use pixel-aligned training examples.
@@
 Conventional feature-based image registration methods are accurate and can model nonlinear deformations but are computationally expensive and sensitive to image degradation.
-As we assume deformations to be approximately linear at a local scale, we take advantage of Fourier-based image correlations Tong et al., (2019), which can efficiently recover a translation between images in the frequency domain.
+As we assume deformations to be approximately linear at a local scale, we take advantage of Fourier-based image correlations tong2019, which can efficiently recover a translation between images in the frequency domain.
 2.3.1 Translational shift
@@
 Both functions $f$ and $g$ represent images of equal height $H$ and width $W$ and are for now assumed to repeat periodically with a periodicity of $H$ and $W$, respectively.
-A common approach to retrieve the translational shift between $f$ and $g$ is to use circular cross-correlation Tong et al., (2019), which can be efficiently computed in the frequency domain as
+A common approach to retrieve the translational shift between $f$ and $g$ is to use circular cross-correlation tong2019, which can be efficiently computed in the frequency domain as
 $\begin{split}\text{CC°}[a,b]&=(f\star g)[a,b]\\ &=\sum_{u,v}f(u-a,v-b)g(u,v)\\ &=\left(\mathcal{F}^{-1}\left\{\overline{\mathcal{F}\{f\}}\mathcal{F}\{g\}\right\}\right)[a,b],\end{split}$ (2)
@@
 $(\Delta u,\Delta v)=\operatorname*{arg\,max}_{(a,b)}\text{CC°}[a,b].$ (3)
-Eq. 2 can be extended to the mean over squared distances between pixel values as
+Eq.˜2 can be extended to the mean over squared distances between pixel values as
 $\begin{split}\text{MSE°}[a,b]&=\frac{1}{HW}\sum_{u,v}(f(u-a,v-b)-g(u,v))^{2}\\ &=\frac{-2(f\star g)[a,b]+\sum_{u,v}(f^{2}(u,v)+g^{2}(u,v))}{HW}.\end{split}$ (4)
-The translational shift $(\Delta u,\Delta v)$ can be recovered analog to Eq.
-3 by computing the $\operatorname*{arg\,min}$.
-For periodic image functions $f$ and $g$, solutions of CC° and MSE° are identical as the sum over squared functions $f^{2}$ and $g^{2}$ is constant Fienup, (1997).
+The translational shift $(\Delta u,\Delta v)$ can be recovered analog to Eq.˜3 by computing the $\operatorname*{arg\,min}$.
+For periodic image functions $f$ and $g$, solutions of CC° and MSE° are identical as the sum over squared functions $f^{2}$ and $g^{2}$ is constant fienup1997.
 Since histological images are not periodic, zero padding is applied to fill both images up to a shape of $(H_{f}+H_{g}-1,W_{f}+W_{g}-1)$, in order to break periodicity and allow processing of images with different heights $H_{f}$, $H_{g}$ and widths $W_{f}$, $W_{g}$.
@@
 Furthermore, additional masks $M_{f}$ and $M_{g}$ are introduced, which have a value of one at all pixel coordinates within original height and width and zero elsewhere.
-We reformulate Eq.
-4 to a non-circular form as
+We reformulate Eq.˜4 to a non-circular form as
 $\text{MSE}[a,b]=\left(\frac{(f_{0}^{2}\star M_{g})-2(f_{0}\star g_{0})+(M_{f}\star g_{0}^{2})}{M_{f}\star M_{g}}\right)[a,b],$ (5)
-which can be efficiently computed by multiple applications of Eq.
-2.
+which can be efficiently computed by multiple applications of Eq.˜2.
 Here, the cross-correlations of $f_{0}^{2}$ and $g_{0}^{2}$ with masks $M_{g}$ and $M_{f}$, respectively, ensure that pixel values of the original unpadded images are not compared with zero padding values.
@@
 2.3.2 Rotation and scale
-While relative scale and rotation between images can also be retrieved in the frequency domain Sheng, (1989); Reddy and Chatterji, (1996), we expect only small relative rotation angles and minor scale variations due to the initial affine registration by matching blood vessels.
+While relative scale and rotation between images can also be retrieved in the frequency domain sheng1989; reddy1996, we expect only small relative rotation angles and minor scale variations due to the initial affine registration by matching blood vessels.
 Therefore, we leverage the parallel processing capability of GPUs to perform an exhaustive search over a fixed set of rotation angles without scaling adjustments.
@@
 Predictions $G(x)$ of the model are compared with actual target images $y$.
-Our style transfer algorithm consists of three main elements:
-  1. 1.
-    Reconstruction Loss $\mathcal{L}_{R}$ : Similar to Isola et al., (2017), we use an $L_{1}$ loss between target $y$ and prediction $G(x)$ to encourage pixel correspondence as
-    $\mathcal{L}_{R}=\mathbb{E}_{x,y}\left\lVert R(y)-G(x)\right\rVert_{1},$ (6)
-    where $R$ performs the proposed online registration to spatially align $y$ and $G(x)$ before loss calculation.
-This enables the utilization of imperfectly aligned training data to penalize any discrepancies in corresponding pixel values between $G(x)$ and $y$.
-  2. 2.
-    Style Loss $\mathcal{L}_{S}$: We apply a texture-sensitive style loss proposed by Gatys et al., (2015) based on squared distances between Gram matrix representations of neural network features.
-Given a pre-trained VGG encoder Simonyan and Zisserman, (2015), Gram matrix representations are computed from feature activations of its layers to characterize the texture of images at different complexities.
+In our case, we translate from the domain of 3D-PLI images to Cresyl violet.
+The translation is performed by a U-Net ronneberger2015 serving as the generator model, which forms the core of our image-to-image translation framework, as illustrated in Fig.˜3.
+The three loss components used to train the U-Net are described below.
+2.4.1 Reconstruction Loss
+Similar to isola2017, we use an $L_{1}$ loss between target $y$ and prediction $G(x)$ to encourage pixel correspondence in the reconstruction loss
+$\mathcal{L}_{R}=\mathbb{E}_{x,y}\left[\left\lVert R_{y,G(x)}(y)-G(x)\right\rVert_{1}\right],$ (6)
+where $R_{y,G(x)}$ performs the proposed online registration to spatially align $y$ and $G(x)$ before loss calculation.
+This enables utilization of imperfectly aligned training data to penalize any discrepancies in corresponding pixel values between $G(x)$ and $y$.
+To disable online registration, $R_{y,G(x)}$ can be replaced by the identity function.
+2.4.2 Style Loss
+We consider two alternative implementations of a style loss $\mathcal{L}_{S}$, both focussing on style preservation. We refer to the first one as Gram loss, and the second as GAN loss.
+For the Gram loss, we apply a texture-sensitive style loss proposed by gatys2015a based on squared distances between Gram matrix representations of neural network features.
+Given a pre-trained VGG encoder simonyan2015, Gram matrix representations are computed from feature activations of its layers to characterize the texture of images at different complexities.
 For each layer $l$, the encoder produces a different number of $N_{l}$ feature maps, each storing $K_{l}$ spatial entries (i.e., height $\times$ width).
-Elements of the Gram matrix $G^{l}_{ij}$ at layer $l$ are computed as the inner product between the $i$-th and $j$-th feature map $F_{i}^{l}$ and $F_{j}^{l}$, where each map is flattened to a $K_{l}$-dimensional vector:
-    $G^{l}_{ij}=\sum^{K_{l}}_{k=1}F^{l}_{ik}F^{l}_{jk}.$ (7)
+Elements of the Gram matrix $\Gamma^{l}_{ij}$ at layer $l$ are computed as the inner product between the $i$-th and $j$-th feature map $F_{i}^{l}$ and $F_{j}^{l}$, where each map is flattened to a $K_{l}$-dimensional vector:
+$\Gamma^{l}_{ij}=\sum^{K_{l}}_{k=1}F^{l}_{ik}F^{l}_{jk}.$ (7)
 Since the Gram matrix computation captures global feature correlations rather than spatial locations, this allows a translation-invariant comparison of image statistics.
-The style loss is computed over all $L$ layers of the VGG encoder, using Gram matrix representations $G^{l}$ for the online registered target and $\hat{G}^{l}$ for the prediction:
-    $\mathcal{L}_{S}=\mathbb{E}_{x,y}\left[\sum^{L}_{l=1}\frac{1}{K_{l}^{2}N_{l}^{2}}\sum^{N_{l}}_{i=1}\sum^{N_{l}}_{j=1}(G^{l}_{ij}-\hat{G}^{l}_{ij})^{2}\right].$ (8)
-    We refer to this style loss as Gram loss.
-  3. 3.
-    Equivariance Loss $\mathcal{L}_{E}$: By registering target $y$ to generator prediction $G(x)$ before loss calculation, displacing objects in $G(x)$ relative to its input $x$ is not penalized by reconstruction loss $\mathcal{L}_{R}$.
-To prevent pixel shifts in $G(x)$, we introduce a third loss component enforcing equivariance with respect to rotations as
-    $\mathcal{L}_{E}=\mathbb{E}_{x}\left\lVert\Omega(G(x))-G(\Omega(x))\right\rVert_{2},$ (9)
-    where operator $\Omega$ represents an image rotation of 180°180°180  °.
-Minimizing Eq.
-9 ensures that pixels in $G(x)$ correspond to pixels at the same pixel coordinates in $x$ as any discrepancy would cause a mismatch of $\Omega(G(x))$ and $G(\Omega(x))$.
-All components are aggregated into total loss
+The style loss is computed over all $L$ layers of the VGG encoder, using Gram matrix representations $\Gamma^{l}$ for the online registered target and $\hat{\Gamma}^{l}$ for the prediction:
+$\mathcal{L}_{S}=\mathbb{E}_{x,y}\left[\sum^{L}_{l=1}\frac{1}{K_{l}^{2}N_{l}^{2}}\sum^{N_{l}}_{i=1}\sum^{N_{l}}_{j=1}(\Gamma^{l}_{ij}-\hat{\Gamma}^{l}_{ij})^{2}\right].$ (8)
+For the GAN loss, we consider adversarial training goodfellow2014 to compare with previous work in virtual staining rivenson2019; cross-zamirski2022.
+We implement the GAN loss in the form of a Wasserstein GAN arjovsky2017a.
+In contrast to conditional GAN training mirza2014; isola2017, we do not condition the discriminator on input images $x$, as this would cause the model to reproduce any misalignment in the training data.
+2.4.3 Equivariance Loss
+By registering target $y$ to generator prediction $G(x)$ before loss calculation, displacements of objects in $G(x)$ relative to $x$ are not captured by $\mathcal{L}_{R}$.
+To prevent pixel shifts in $G(x)$, we enforce equivariance with respect to rotations through the equivariance loss
+$\mathcal{L}_{E}=\mathbb{E}_{x}\left[\left\lVert\Omega(G(x))-G(\Omega(x))\right\rVert_{2}\right],$ (9)
+where operator $\Omega$ represents an image rotation of 180°.
+Minimizing Eq.˜9 ensures that pixels in $G(x)$ correspond to pixels at the same pixel coordinates in $x$ as any discrepancy would cause a mismatch of $\Omega(G(x))$ and $G(\Omega(x))$.
+2.4.4 Total loss formulation
+All components are aggregated into the total loss
 $\mathcal{L}=\lambda\mathcal{L}_{R}+(1-\lambda)\mathcal{L}_{S}+\eta\mathcal{L}_{E},$ (10)
-with relative weightings $\lambda\in[0,1]$ and $\eta\geq 0$ as hyperparameters. An overview of the framework is illustrated in Fig. 3.
-As an alternative for style loss $\mathcal{L}_{S}$, we consider adversarial training Goodfellow et al., (2014) using a GAN loss to compare with previous work in virtual staining Rivenson et al., (2019); Cross-Zamirski et al., (2022).
-We implement the GAN loss in the form of a Wasserstein GAN Arjovsky et al., (2017).
-In contrast to conditional GAN training Mirza and Osindero, (2014); Isola et al., (2017), we do not condition the discriminator on input images $x$, as this would cause the model to reproduce any misalignment in the training data.
-We refer to models trained with Gram loss for $\mathcal{L}_{S}$ as Gram and models trained with GAN loss as GAN.
-In combination with the online registration head, we refer to the models as Gram+Reg and GAN+Reg.
+with relative weightings $\lambda\in[0,1]$ and $\eta\geq 0$ as hyperparameters.
+We denote models that use Gram loss as the style loss $\mathcal{L}_{S}$ as Gram, and models that use GAN loss as the style loss $\mathcal{L}_{S}$ as GAN.
+When online registration is enabled during computation of reconstruction loss $\mathcal{L}_{R}$, the corresponding models are denoted as Gram+Reg and GAN+Reg.
+Base models Gram and GAN compute the reconstruction loss with online registration disabled.
 2.5 Model training
-For the generator $G$, we use the same 5-layer U-Net Ronneberger et al., (2015) with numbers of features [32, 64, 128, 256, 512] in all experiments, and adjust the input and output channels to 3 according to our setup.
+For the generator $G$, we use the same 5-layer U-Net ronneberger2015 with numbers of features [32, 64, 128, 256, 512] in all experiments, and adjust the input and output channels to 3 according to our setup.
 To train $G$, we use square 3D-PLI patches of 444 pixels size, represented by parameter maps transmittance ($I_{T}$), retardation ($\sin\delta$) and direction ($\varphi$).
@@
 We normalize image pixel values of the Cresyl violet staining to the range of [0, 1].
-For computing style loss $\mathcal{L}_{S}$, we extract features from a VGG19 model Simonyan and Zisserman, (2015) to compute the Gram loss.
-The VGG feature encoder network has a depth of four layers and three input channels with pre-trained weights on ImageNet Deng et al., (2009).
+For computing style loss $\mathcal{L}_{S}$, we extract features from a VGG19 model simonyan2015 to compute the Gram loss.
+The VGG feature encoder network has a depth of four layers and three input channels with pre-trained weights on ImageNet deng2009.
 We multiply the style loss $\mathcal{L}_{S}$ by a constant factor of 104 to bring it to the same order of magnitude as reconstruction loss $\mathcal{L}_{R}$.
-For the training, we use Adam optimizer Kingma and Ba, (2017) with $\beta_{1}$ = 0.9, $\beta_{2}$ = 0.999 and a learning rate of 10-3.
-If not stated otherwise, we use $\eta$ = 0.1 and $\lambda$ = 0.5 as default in the training objective in Eq.
-10.
-In the case of GAN style loss $\mathcal{L}_{S}$, we use a 4-layer convolutional network as discriminator with kernel size 4, stride 2, padding 1, and feature size of [32, 64, 128, 256], followed by batch normalization Ioffe and Szegedy, (2015) and Leaky ReLU after each convolution.
-For training, we use the Wasserstein GAN Arjovsky et al., (2017) objective and a separate Adam optimizer for the discriminator and the generator, using $\beta_{1}$ = 0.5, $\beta_{2}$ = 0.999 and a learning rate of 10-4.
+For the training, we use Adam optimizer kingma2017 with $\beta_{1}$ = 0.9, $\beta_{2}$ = 0.999 and a learning rate of 10-3.
+If not stated otherwise, we use $\eta$ = 0.1 and $\lambda$ = 0.5 as default in the training objective in Eq.˜10.
+In the case of GAN style loss $\mathcal{L}_{S}$, we use a 4-layer convolutional network as discriminator with kernel size 4, stride 2, padding 1, and feature size of [32, 64, 128, 256], followed by batch normalization ioffe2015 and Leaky ReLU after each convolution.
+For training, we use the Wasserstein GAN arjovsky2017a objective and a separate Adam optimizer for the discriminator and the generator, using $\beta_{1}$ = 0.5, $\beta_{2}$ = 0.999 and a learning rate of 10-4.
 We perform five updates for the discriminator for one update of the generator and clamp discriminator weights at 0.03 after each step.
@@
 For model training and evaluation, we use eight coronal sections at the level of the central sulcus.
-Seven sections are used for training and one section is kept for testing with a gap of 0.6 mm between train and test sections (Fig.
-4 A).
-From the training sections, 27 affine-aligned ROIs are extracted (Fig.
-4 B), where one ROI is held out for validation to identify possible overfitting.
+Seven sections are used for training and one section is kept for testing with a gap of 0.6 mm between train and test sections (Fig.˜4A).
+From the training sections, 27 affine-aligned ROIs are extracted (Fig.˜4B), where one ROI is held out for validation to identify possible overfitting.
 For each ROI, we extract joint target Cresyl violet images and 3D-PLI modalities at the same center location.
@@
 Data augmentation
-To enhance the robustness of our trained models we employ 3D-PLI-specific data augmentations, which were carefully modified to keep physically plausible signal parameters Oberstrass et al., (2024).
+To enhance the robustness of our trained models we employ 3D-PLI-specific data augmentations, which were carefully modified to keep physically plausible signal parameters oberstrass2024a.
 Specifically, we perform random rotation by angles between -180° and 180° with mirror padding and horizontal and vertical random flipping.
@@
 Implementation
-All models were trained on the supercomputer JURECA-DC at the Jülich Supercomputing Centre (JSC, Thörnig, 2021) on a single node by splitting each batch equally onto 4 NVIDIA A100 GPUs using distributed data-parallel strategy.
+All models were trained on the supercomputer JURECA-DC at the Jülich Supercomputing Centre (JSC, thornig2021) on a single node by splitting each batch equally onto 4 NVIDIA A100 GPUs using distributed data-parallel strategy.
 For data pre-processing 128 worker processes were spawned on 128 CPU cores.
 For reference, training for 100 epochs took 8 hours with online registration and 4 hours without on this hardware.
-The implementations are based on the Quicksetup-ai template by the HelmholtzAI Consultants Munich [^1] https://github.com/HelmholtzAI-Consultants-Munich/Quicksetup-ai, using the frameworks: PyTorch Paszke et al., (2019), PyTorch Lightning Borovec et al., (2022) and Hydra Yadan, (2019).
+The implementations are based on the Quicksetup-ai template by the HelmholtzAI Consultants Munich [^1] https://github.com/HelmholtzAI-Consultants-Munich/Quicksetup-ai, using the frameworks: PyTorch NEURIPS2019_9015, PyTorch Lightning jirka_borovec_2022_7447212 and Hydra Yadan2019Hydra.
 Online registration
-For the online registration head, we restrict accepted solutions of Eq.
-5 to translation + rotation pairs that cause the registered target to have full overlap with the prediction, avoiding loss calculation over zero-padded values.
+For the online registration head, we restrict accepted solutions of Eq.˜5 to translation + rotation pairs that cause the registered target to have full overlap with the prediction, avoiding loss calculation over zero-padded values.
 We check the translation correction for 31 rotation angles from -7.5° to 7.5° with steps of 0.5° and take the translation + rotation pair with the best registration score.
@@
 3.1.1 Test data
-We manually select four ROIs for model evaluation, ensuring a diverse representation of different cytoarchitectonic characteristics from the held-out test section (Fig.
-4B).
-The ROIs contain primary motor cortex 4a, temporal cortical area TE, the hippocampal cornu Ammonis (CA) region, and parts of the putamen and globus pallidus (Pars interna and Pars externa) as subcortical structures (Fig.
-5).
-As the computation of image metrics requires a precise alignment of test data, we perform elastic registration of test ROIs based on landmarks and image intensity using the bUnwarpJ Arganda-Carreras et al., (2006) algorithm.
+We manually select four ROIs for model evaluation, ensuring a diverse representation of different cytoarchitectonic characteristics from the held-out test section (Fig.˜4B).
+The ROIs contain primary motor cortex 4a, temporal cortical area TE, the hippocampal cornu Ammonis (CA) region, and parts of the putamen and globus pallidus (Pars interna and Pars externa) as subcortical structures (Fig.˜5).
+As the computation of image metrics requires a precise alignment of test data, we perform elastic registration of test ROIs based on landmarks and image intensity using the bUnwarpJ arganda-carreras2006 algorithm.
 We use predictions of an independently trained Gram+Reg model as target.
@@
 Embedded windows show magnified details inside each ROI.
-The locations of the magnified views are indicated by yellow frames.
 Columns each show one of the four test ROIs taken from the anterior subdivision of the primary motor cortex (4a), the hippocampal cornu Ammonis (CA) region, temporal cortical area TE, and parts of the putamen and globus pallidus (GP; a: GP Pars interna; b: GP Pars externa) as subcortical nuclei.
-The first three rows demonstrate 3D-PLI modalities transmittance $I_{T}$, retardation $\sin\delta$ (scaled using gamma correction for visualization) and fiber orientation in HSV color space (hue: direction $\varphi$; saturation, brightness: retardation $\sin\delta$).
+The first three rows demonstrate 3D-PLI modalities transmittance, retardation (scaled using gamma correction for visualization) and fiber orientation in HSV color space (hue: fiber direction; saturation/brightness: retardation).
 The 3D-PLI modalities are compared to the registered target Cresyl violet and predicted virtual staining.
 3.1.2 Evaluation scores
-We evaluate the impact of different model parameter choices on the quality of the predicted virtual staining by applying structural similarity index measure (SSIM, Wang et al., 2004), mutual information (MI), and root-mean-square error (RMSE).
+We evaluate the impact of different model parameter choices on the quality of the predicted virtual staining by applying structural similarity index measure (SSIM, wang2004), mutual information (MI), and root-mean-square error (RMSE).
 For each metric, we report the mean over all test ROIs.
-To evaluate the ability to localize cell bodies, we compute F1 scores for instance segmentations of predicted and target cells using a contour proposal network (CPN, Upschulte et al., 2022, 2023).
-We only consider cells with a minimum in-plane size of 100 µm² and require a minimum intersection over union (IoU) of 30%.
-Smaller cells could not be reconstructed accurately and would introduce noise into the evaluation.
-For the cell detection model, we fine-tuned a pre-trained222ginoro_CpnResNeXt101UNet-fbe875f1a3e5ce2c CPN for cell body segmentation in cell-stained microscopy images using the celldetection [^3] https://github.com/FZJ-INM1-BDA/celldetection Python package.
+To evaluate how well cell positions are preserved by different models, we compute F1 scores based on cell instance segmentations by a contour proposal network (CPN, upschulte2022; upschulte2023).
+Predicted cell instances are obtained from segmentations of the virtual staining and compared to target cell instances from the corresponding Cresyl violet images.
+For each ROI, predicted and target cells are matched by calculating their intersection over union (IoU), requiring a minimum IoU of 30% for a match.
+Each target cell can be a match for at most one predicted cell.
+Matched instances are counted as true positives, unmatched predicted cells as false positives, and unmatched target cells as false negatives.
+F1 scores are then computed from aggregated counts across all ROIs.
+For the cell detection model, we fine-tuned a pre-trained CPN222ginoro_CpnResNeXt101UNet-fbe875f1a3e5ce2c for cell body segmentation in cell-stained microscopy images using the celldetection [^3] https://github.com/FZJ-INM1-BDA/celldetection Python package.
 Fine-tuning was performed on a diverse mix of manually annotated images, as well as synthetic data.
@@
 Figure 6: Robustness of different online registration metrics against synthetic image distortions.
-We report the proportion of correct rigid alignments for different blur (A) and noise (B) levels by registration metrics MSE, CC, PC, and BIPC.
+We report the proportion of correct rigid alignments for different blur (A) and noise (B) levels by registration metrics MSE, $C$, PC, and BIPC.
 The registration is considered successful (a hit) if the translational displacement obtained remains within 5 pixels of the actual displacement.
-To compute a pixel-aligned reconstruction loss we apply the online registration head (Sec.
-2.3).
-We consider correlation-based registration metrics CC and MSE with phase correlation (PC, Kuglin, 1975) and blur-invariant phase correlation (BIPC, Ojansivu and Heikkila, 2007).
-We apply a Hann window to PC and BIPC before their calculation to mitigate their bias towards the sharp image edges Gonzales and Woods, (2008).
+To compute a pixel-aligned reconstruction loss we apply the online registration head (Sec.˜2.3).
+We consider correlation-based registration metrics CC and MSE with phase correlation (PC, kuglin1975) and blur-invariant phase correlation (BIPC, ojansivu2007).
+We apply a Hann window to PC and BIPC before their calculation to mitigate their bias towards the sharp image edges gonzalez2008digital.
 To understand how image degradation effects can influence the success of the online registration head, we compare the robustness of the metrics against noise and blur.
@@
 The registration is considered successful if the determined translational displacement remains within 5 pixels of the actual displacement.
-Fig.
-6 shows that as blur increases, BIPC with a Hann window achieves the best performance.
+Fig.˜6 shows that as blur increases, BIPC with a Hann window achieves the best performance.
 This result is expected, as the metric remains invariant to blur.
@@
 Table 1: Performance of models trained with different online registration metrics in terms of similarity of generated images with ground truth.
-We compare cross-correlation (CC), phase-correlation (PC), blur-invariant phase-correlation(BIPC), and mean-squared error (MSE) as online registration metrics against using no online registration (-).
+We compare cross-correlation ($C$), phase-correlation (PC), blur-invariant phase-correlation(BIPC), and mean-squared error (MSE) as online registration metrics against using no online registration (-).
+Arrows indicate the direction of better performance ($\uparrow$ higher is better, $\downarrow$ lower is better).
+Best scores per column in bold.
 Method Metric MI $\uparrow$ RMSE $\downarrow$ SSIM $\uparrow$ F1 $\uparrow$
-Gram - 0.126 33.6 0.354 48.8
-Gram+Reg CC 0.136 33.0 0.379 55.2
-BIPC 0.149 31.5 0.371 54.0
-PC 0.185 31.9 0.415 59.4
-MSE 0.226 29.8 0.444 63.7
+Gram - 0.126 33.6 0.354 30.2
+Gram+Reg $C$ 0.136 33.0 0.379 34.4
+BIPC 0.149 31.5 0.371 32.6
+PC 0.185 31.9 0.415 38.0
+MSE 0.226 29.8 0.444 41.3
 To make an optimal choice of a registration metric for online registration, we train Gram+Reg models with CC, BIPC, PC, and MSE compared to a Gram model without online registration.
@@
 For Gram+Reg models trained with CC, BIPC and the Gram model, $\lambda=0.1$ was chosen as it improved results compared to $\lambda=0.5$ for PC and MSE.
-Tab.
-1 shows a quantitative comparison of the models.
+Tab.˜1 shows a quantitative comparison of the models.
 Independent of the metric, the best results are observed when using Gram+Reg with online registration Among all metrics considered, the model with MSE performs best.
 3.3 Performance analysis of different model variants
-Figure 7: Including equivariance loss $\mathcal{L}_{E}$ in the training improves cell instance overlap of Gram+Reg and GAN+Reg models.
-(A) 3D-PLI input as transmittance $I_{T}$, retardation $\sin\delta$ (scaled using gamma correction) and fiber orientation map in HSV color space (hue: direction $\varphi$; saturation, brightness: retardation $\sin\delta$) for an example patch from the pyramidal layer of the hippocampal CA1 region.
+Figure 7: Example image illustrating how equivariance loss $\mathcal{L}_{E}$ improves cell instance overlap of Gram+Reg and GAN+Reg models.
+(A) 3D-PLI input as transmittance, retardation (scaled using gamma correction) and fiber orientation map in HSV color space (hue: fiber direction; saturation/brightness: retardation).
 (B) Registered target Cresyl violet staining, a cell segmentation by a CPN model, and the identified target cell instances.
 (C) Predicted virtual staining of models trained with equivariance loss $\mathcal{L}_{E}$ enabled (✓) or disabled (✗).
-Cell instances are overlaid with their target cells.
-Models trained with equivariance loss show less displacement of cells and increased F1 scores for this specific image patch.
-Since the displacement occurs randomly, a patch and random seeds for the models are selected that show a particularly strong effect.
-To justify including equivariance loss $\mathcal{L}_{E}$ (Eq.
-9) in the total loss, Fig.
-7 illustrates the effect of training Gram+Reg models with this loss component enabled and disabled.
+Predicted cell instances (red) by each model are overlaid with the target cells (gray) and F1 scores computed for cells in this specific image patch.
+Since the displacement occurs randomly, a patch from our test set and random seeds for the models were manually selected that illustrate that effect.
+The patch shows the pyramidal layer of the hippocampal CA1 region.
+To justify including equivariance loss $\mathcal{L}_{E}$ (Eq.˜9) in the total loss, Fig.˜7 illustrates the effect of training different models with this loss component enabled and disabled.
 With $\mathcal{L}_{E}$ disabled ($\eta$ = 0), cell instances in the model predictions exhibit a relative displacement.
@@
 The deviation is reported as standard error over four independent trainings with different random seeds.
+Arrows indicate the direction of better performance ($\uparrow$ higher is better, $\downarrow$ lower is better).
+Best scores per column in bold.
 $\mathcal{L}_{S}$ Reg. $\mathcal{L}_{E}$ MI $\uparrow$ RMSE $\downarrow$ SSIM $\uparrow$ F1 $\uparrow$
-Gram 0.224 $\pm$ 0.003 29.6 $\pm$ 0.6 0.445 $\pm$ 0.003 63.1 $\pm$ 0.5
-0.211 $\pm$ 0.004 30.2 $\pm$ 0.4 0.432 $\pm$ 0.004 61.6 $\pm$ 0.4
-0.111 $\pm$ 0.012 33.6 $\pm$ 0.8 0.314 $\pm$ 0.023 38.9 $\pm$ 7.6
-0.107 $\pm$ 0.007 34.2 $\pm$ 0.3 0.304 $\pm$ 0.015 40.6 $\pm$ 5.7
-GAN 0.160 $\pm$ 0.008 30.5 $\pm$ 0.5 0.389 $\pm$ 0.012 50.3 $\pm$ 2.1
-0.126 $\pm$ 0.008 33.0 $\pm$ 0.4 0.347 $\pm$ 0.010 43.0 $\pm$ 1.4
-0.085 $\pm$ 0.008 38.1 $\pm$ 0.5 0.249 $\pm$ 0.001 21.1 $\pm$ 0.3
-0.096 $\pm$ 0.008 38.8 $\pm$ 1.7 0.253 $\pm$ 0.005 22.2 $\pm$ 1.3
-To quantify the impact of the loss component $\mathcal{L}_{E}$ on the predictions, we compare the models with $\mathcal{L}_{E}$ enabled and disabled in Tab.
-2.
+Gram 0.224 $\pm$ 0.002 29.6 $\pm$ 0.4 0.445 $\pm$ 0.002 41.0 $\pm$ 0.4
+0.211 $\pm$ 0.004 30.2 $\pm$ 0.4 0.432 $\pm$ 0.004 35.2 $\pm$ 1.2
+0.111 $\pm$ 0.012 33.6 $\pm$ 0.8 0.314 $\pm$ 0.023 22.6 $\pm$ 4.8
+0.107 $\pm$ 0.007 34.2 $\pm$ 0.3 0.304 $\pm$ 0.015 22.5 $\pm$ 3.1
+GAN 0.160 $\pm$ 0.008 30.5 $\pm$ 0.5 0.389 $\pm$ 0.012 34.7 $\pm$ 1.6
+0.126 $\pm$ 0.008 33.0 $\pm$ 0.4 0.347 $\pm$ 0.010 28.0 $\pm$ 1.3
+0.085 $\pm$ 0.008 38.1 $\pm$ 0.5 0.249 $\pm$ 0.001 12.9 $\pm$ 0.3
+0.096 $\pm$ 0.008 38.8 $\pm$ 1.7 0.253 $\pm$ 0.005 13.9 $\pm$ 1.5
+To quantify the impact of the loss component $\mathcal{L}_{E}$ on the predictions, we compare the models with $\mathcal{L}_{E}$ enabled and disabled in Tab.˜2.
 We consider Gram and Gram+Reg, as well as GAN and GAN+Reg as model variants.
@@
 Without online registration, the advantage of including $\mathcal{L}_{E}$ becomes negligible.
-Furthermore, results in Tab.
-2 demonstrate that Gram models outperform their corresponding GAN models in all evaluation metrics.
+Furthermore, results in Tab.˜2 demonstrate that Gram models outperform their corresponding GAN models in all evaluation metrics.
 Regardless of the variant used for style loss $\mathcal{L}_{S}$, all models benefit from the online registration.
+We present averaged scores across all test ROIs.
+Individual scores per ROI, are provided in Tab.˜5 in the appendix as additional information.
 Figure 8: Comparison of virtual stainings for a patch of temporal area TE.
-(A) 3D-PLI input visualized as fiber orientation in HSV color space (hue: direction $\varphi$; saturation, brightness: retardation $\sin\delta$) and transmittance $I_{T}$.
+(A) 3D-PLI input visualized as fiber orientation in HSV color space (hue: fiber direction; saturation/brightness: retardation) and transmittance.
 (B) Registered target Cresyl violet staining for reference with annotations of cortical layers I-VI and white matter (wm).
@@
 (D) GLI images ranging from the pial surface to the cortex/white matter transition and average profiles (red lines) for layers II - VI.
-A comparison of model predictions is shown in Fig.
-8C for the whole cortical depth of an Isocortex sample from temporal cortical area TE.
+A comparison of model predictions is shown in Fig.˜8C for the whole cortical depth of an Isocortex sample from temporal cortical area TE.
 GAN+Reg and Gram+Reg provide the most realistic-looking reconstructions in terms of relative size and shape of cell bodies, along with the differences between layers concerning cell packing density.
@@
 Both methods introduce an artificial arrangement of cells into cortical columns not present in the original Cresyl violet staining.
-To further analyze the ability of the methods to reconstruct the laminar cell organization, we compute grey level index (GLI) values Schleicher et al., (1978); Zilles et al., (1978).
-GLI values provide an established proxy for volume density of stained cell bodies in gray matter regions and are used to characterize the laminar architecture of cortical areas Schleicher et al., (2000).
+To further analyze the ability of the methods to reconstruct the laminar cell organization, we compute grey level index (GLI) values zilles1978.
+GLI values provide an established proxy for volume density of stained cell bodies in gray matter regions and are used to characterize the laminar architecture of cortical areas schleicher2000.
 To compute GLI values, adaptive thresholding is applied to create a mask of pixels occupied by cell bodies for each image.
 The masks are subsequently down scaled by a factor of 16 to a spatial resolution of 20.8 µm, with each value representing the fraction of segmented pixels.
-Resulting GLI images, cropped to the area between pial surface and the cortex/white matter transition, are displayed for the Cresyl violet target and the predictions in Fig.
-8D.
+Resulting GLI images, cropped to the area between pial surface and the cortex/white matter transition, are displayed for the Cresyl violet target and the predictions in Fig.˜8D.
 For each of the GLI images, 31 intensity profiles vertically oriented to the cortical layers are extracted to represent the columnar distribution of GLI values.
@@
 To obtain one representative profile for each image, the profiles are averaged and smoothed using a mean filter with kernel size 3 to improve the signal-to-noise ratio.
-The average profiles are shown on top of each GLI image in Fig.
-8D.
+The average profiles are shown on top of each GLI image in Fig.˜8D.
 The average GLI profile for the GAN model shows no clear distinction between cortical layers II-IV and misses a peak for layer IV.
@@
 Gram+Reg shows the clearest distinction between layers and even replicates slightly nuanced peaks of the profile within layers.
-3.4 Optimal balancing of style and content
+3.4 Effect of loss weighting parameters
+Table 3: Ablation study on weighting parameter $\eta$ of equivariance loss $\mathcal{L}_{E}$ for a Gram+Reg model.
+Performance is robust across a wide range of $\eta>0$ values.
+Only when the loss term is completely removed ($\eta=0$), we observe a measurable degradation in performance.
+The deviation is reported as standard error over four independent trainings with different random seeds.
+Arrows indicate the direction of better performance ($\uparrow$ higher is better, $\downarrow$ lower is better).
+Best scores per column in bold.
+Method $\eta$ MI $\uparrow$ RMSE $\downarrow$ SSIM $\uparrow$ F1 $\uparrow$
+Gram+Reg 0.0 0.211 $\pm$ 0.004 30.2 $\pm$ 0.4 0.432 $\pm$ 0.004 35.2 $\pm$ 1.2
+0.01 0.225 $\pm$ 0.002 29.5 $\pm$ 0.3 0.446 $\pm$ 0.001 41.6 $\pm$ 0.2
+0.1 0.224 $\pm$ 0.003 29.6 $\pm$ 0.6 0.445 $\pm$ 0.003 41.0 $\pm$ 0.4
+1.0 0.220 $\pm$ 0.001 29.4 $\pm$ 0.1 0.445 $\pm$ 0.001 41.4 $\pm$ 0.2
+10.0 0.224 $\pm$ 0.002 29.4 $\pm$ 0.3 0.446 $\pm$ 0.001 41.2 $\pm$ 0.2
+To incorporate equivariance loss $\mathcal{L}_{E}$ into previous models, we set its weighting parameter to $\eta=0.1$.
+To validate this choice, we perform training of Gram+Reg models with $\eta$ values ranging from 0 to 10.
+Results in Tab.˜3 indicate that performance is robust for $\eta>0$.
+Only when the equivariance loss is entirely removed from training (setting $\eta=0$), we observe a measurable negative effect on the results.
+We report mean values and standard errors over four independent trainings to show the significance of this effect.
 Table 4: Different values for weighting parameter $\lambda$ show that a balance between style loss $\mathcal{L}_{S}$ and reconstruction loss $\mathcal{L}_{R}$ is required in the training of a Gram+Reg model.
@@
 Setting $\lambda=1$ means only $\mathcal{L}_{R}$ is used.
+Arrows indicate the direction of better performance ($\uparrow$ higher is better, $\downarrow$ lower is better).
+Best scores per column in bold.
 Method $\lambda$ MI $\uparrow$ RMSE $\downarrow$ SSIM $\uparrow$ F1 $\uparrow$
-Gram+Reg 0 0.124 33.8 0.340 48.7
-0.03 0.162 31.9 0.394 57.7
-0.25 0.210 30.7 0.436 62.8
-0.5 0.226 29.8 0.444 63.7
-0.75 0.238 29.8 0.448 63.2
-0.97 0.230 30.7 0.446 59.6
-1.0 0.090 38.0 0.347 0.0
+Gram+Reg 0 0.124 33.8 0.340 28.6
+0.03 0.162 31.9 0.394 36.1
+0.25 0.210 30.7 0.436 39.8
+0.5 0.226 29.8 0.444 41.3
+0.75 0.238 29.8 0.448 41.4
+0.97 0.230 30.7 0.446 40.2
+1.0 0.090 38.0 0.347 0
 Figure 9: Predicted virtual Cresyl violet stainings for different weightings between reconstruction and style loss components.
-(A) 3D-PLI inputs as transmittance $I_{T}$, retardation $\sin\delta$ and fiber orientation in HSV color space (hue: direction $\varphi$; saturation, brightness: retardation $\sin\delta$).
+(A) 3D-PLI inputs as transmittance, retardation and fiber orientation in HSV color space (hue: fiber direction; saturation/brightness: retardation).
 The retardation has been scaled using gamma correction for visualization.
@@
 (iv) Layer V of the primary motor cortex with arrowheads highlighting two Betz cells.
-To identify an optimal weighting $\lambda$ of the style loss in Eq.
-10, we train Gram+Reg models with different values for $\lambda$ and MSE online registration metric.
-A quantitative evaluation of the trained models in Tab.
-3 shows that larger values for $\lambda$ up to $\lambda$ = 0.75 achieve the best MI, RMSE, and SSIM scores.
-This indicates that the reconstruction loss should be given a stronger weight but should not be used alone ($\lambda=1$).
-The highest overlap of predicted cell instances measured by the F1 score is obtained for a balanced weighting with $\lambda$ = 0.5.
-A visual comparison of predictions of Gram+Reg models trained with different choices for $\lambda$ is shown in Fig.
-9 for selected crops from the test data.
+To identify an optimal weighting $\lambda$ of the style loss in Eq.˜10, we train Gram+Reg models with different values for $\lambda$ and MSE online registration metric.
+A quantitative evaluation of the trained models in Tab.˜4 shows that larger values for $\lambda$ up to $\lambda$ = 0.75 achieve the best evaluation scores.
+This indicates that the reconstruction loss should be given a stronger weight but should not be used exclusively ($\lambda=1$).
+A visual comparison of predictions of Gram+Reg models trained with different choices for $\lambda$ is shown in Fig.˜9 for selected crops from the test data.
 By using $\mathcal{L_{R}}$ only (i.e.
@@
 However, generators trained with an overweighting of style loss also tend to miss some strongly pronounced structures in the predictions, such as the blood vessels shown in (i) or the Betz cells in (iv).
-3.5 Reliability of predicted cell instances across cell sizes
-Figure 10: Reliability of predicted cell instances increases with their size.
+3.5 Influencing factors on model predictions
+To assess the reliability of our method under varying biological and imaging conditions, we analyze in how far the reconstruction of cell instances depends on cell size and local strength of birefringence.
+We also examine how different focus levels of the LMP-1 impair the quality of the virtual staining.
+Figure 10: Reliability of predicted cell instances increases with their size and decreases with a stronger retardation.
 We report F1 scores between detected cell instances by a CPN model in the predicted and target stainings.
-F1 scores are reported for intervals of the segmented in-plane cell body area.
+F1 scores are computed for intervals of (A) the segmented in-plane cell body area and (B) smoothed retardation maps.
+Retardation values are sampled at each cell location.
 Each bar represents the average over four independently trained models.
@@
 To count false negatives, target cells within that range are matched with all cells in the prediction.
-Cell sizes are corrected with the 2D shrinkage factor of 0.97 estimated in Sec.
-2.1.2, assuming homogeneous shrinkage of cells and surrounding tissue.
-Results in Fig.
-10 show that smaller cells < 50 µm have much lower F1 scores, below 20.0, with all methods.
+Results in Fig.˜10A show that reconstruction of smaller cells < 50 µm² has much lower F1 scores, below 20.0, throughout all methods.
 With increasing 2D cell size, they can be identified more accurately.
 The Gram+Reg model shows highest F1 scores across all cell sizes and smallest variation between four independently trained models, measured as standard deviation.
+Birefringent tissue, such as myelinated nerve fibers, can obscure signals from other components in 3D-PLI.
+To examine whether this effect impacts the prediction of cell instances in the virtual staining, we use retardation maps as a measure of birefringence strength.
+The maps are smoothed with a 10 pixels square median kernel to reduce local variance and obtain values representative of the surrounding tissue area.
+For each cell instance segmented by the CPN model in the virtual staining, we sample a retardation value at its center location.
+Values are grouped into intervals from 0.0 to 0.3 in steps of 0.05.
+Since we focus the analysis on gray matter, larger retardation values do not occur.
+As shown in Fig.˜10B, the ability of all models to reconstruct cell instances decreases significantly for cells dominated by stronger retardation signals.
+For Gram+Reg this effect is minimal, performing overall best.
+Figure 11: Virtual staining quality varies with the focus level of the LMP-1 microscope, which influences the captured diffraction patterns.
+Example images show the vervet entorhinal cortex with zoom-ins to the architecture of cortical layers II-III.
+3D-PLI inputs are represented as transmittance, retardation (scaled with gamma correction) and fiber orientation in HSV color space (hue: fiber direction; saturation/brightness: retardation).
+The same Gram+Reg model is used to predict virtual stainings from a focus level adjusted to the tissue center (+0 µm) and 30 µm above and below.
+We observe that the virtual staining intensities are not always homogenous within generated images.
+Such effects can especially be observed as cloud-like patterns in Fig.˜5 for the predicted virtual staining of ROIs showing the primary motor cortex and subcortical nuclei.
+These variations coincide with variations in the diffraction pattern of the associated transmittance maps, which in turn can be influenced by the focus level of the LMP-1 microscope.
+To better quantify this relationship, we apply a Gram+Reg model to predict virtual stainings from an example image showing the vervet entorhinal cortex, captured at three focus levels: centered on the tissue (+0 µm), -30 µm and +30 µm.
+Results are summarized in Fig.˜11.
+We observe the highest consistency in staining intensity at the lowest focus level of (-30 µm).
+Higher focus levels introduce an increasing amount of staining artifacts.
 4 Discussion
+Our ablation experiments in Sec.˜3.2, Sec.˜3.3 and Sec.˜3.4 support the design decisions of the proposed virtual staining method.
+In Sec.˜3.5 we investigated several influencing factors on the predictions.
+In the following, we want to discuss the design decisions of the proposed virtual staining method and assess its reliability for downstream analysis.
 We introduced an online registration head capable of approximating smooth nonlinear deformations, given sufficiently small patch sizes for training.
@@
 We identified MSE in the Fourier domain through multiple applications of cross-correlation as superior to other typically used metrics in Fourier-based image registration, including PC.
-This is a notable observation, as PC is a popular choice due to its robustness to intensity variations and frequency-dependent noise Tong et al., (2019).
+This is a notable observation, as PC is a popular choice due to its robustness to intensity variations and frequency-dependent noise tong2019.
 However, during early training stages, model predictions are often blurred.
@@
 Additionally, random weight initialization introduces noise in the first steps.
-MSE appears more resilient to such blur and frequency-independent noise, leading to more accurate predictions (Fig.
-6).
-This behavior can be explained by PC’s reliance on strong high-frequency spectral peaks, which are diminished by blur Pedone et al., (2013).
+MSE appears more resilient to such blur and frequency-independent noise, leading to more accurate predictions (Fig.˜6).
+This behavior can be explained by PC’s reliance on strong high-frequency spectral peaks, which are diminished by blur pedone2013.
 MSE, in contrast, maintains sensitivity across the frequency spectrum and can better handle smooth, low-detail inputs that dominate during initial training.
@@
 In other words, there is a cross-dependency of accurate pixel reconstructions and producing a high contrast over training iterations, which may result in a conflict during training.
-This can be seen in Fig.
-9, where a Gram+Reg model focussing on pixel reconstruction only made heavily blurred predictions that do not provide structural details for cell-precise registration.
-Adding a style loss component leads to more pronounced contours while simultaneously achieving higher pixel reconstruction accuracy (Tab.
-3) through online registration.
+This can be seen in Fig.˜9, where a Gram+Reg model focussing on pixel reconstruction only made heavily blurred predictions that do not provide structural details for cell-precise registration.
+Adding a style loss component leads to more pronounced contours while simultaneously achieving higher pixel reconstruction accuracy (Tab.˜4) through online registration.
 This underpins the importance of the additional style loss, which does not depend on perfect alignment and guides the training to produce structures.
+Still, some manual pre-alignment was required to keep the registration error of training data within 70 pixels (91 µm) and ensure sufficient structural overlap between training patches.
+This limited the amount of available training data, as larger registration errors, such as those from the initial affine alignment of whole sections, were too large for the online registration to succeed.
 Apparently, the choice for Gram or GAN as style loss should be considered in the light of the application.
-Training with Gram loss led to more accurate cell localization and pixel values (Tab.
-2).
-However, it was affected by staining intensity inhomogeneities, resulting in blurry cells (Fig.
-8).
+Training with Gram loss led to more accurate cell localization and pixel values (Tab.˜2).
+However, it was affected by staining intensity inhomogeneities, resulting in blurry cells (Fig.˜8).
 GAN loss, on the other hand, produced higher contrast in cells but with lower accuracy.
@@
 It requires no balancing of model capacities, is robust to loss balancing, and does not collapse during training.
+We expect that the expressiveness of the Gram matrix representation could be enhanced by replacing the VGG encoder, pre-trained on ImageNet, with a domain-specific model, which is powerful on microscopy data.
+This is technically motivated by the observation that higher-level features in ImageNet-pretrained networks, such as VGG, are optimized for object semantics in natural images (e.g., faces, animals, vehicles), which are irrelevant in the context of fine-grained texture characteristic of microscopy images stuckner2022.
+In contrast, microscopy-specific encoders are better suited to capture such low- to mid-level textural features and have been shown to improve clustering oberstrass2024a, segmentation stuckner2022, and the evaluation of generative models kropp2024 over those trained on ImageNet.
+Therefore, computation of Gram matrices using domain-specific encoders spitzer2017; schiffer2021; schiffer2021a or emerging foundation models for histology tran2024 might further enhance representation of cytoarchitectural characteristics.
 Throughout all models, we observed occasional staining inhomogeneities in the predicted virtual stainings.
-This effect was particularly pronounced in the form of stripe patterns in the primary motor cortex ROI (Fig.
-5).
+An example of local-scale variation is the inhomogeneity observed in Fig.˜8, which may be attributed to the tiling strategy used during model inference to transfer whole sections.
+Since the model lacks a global view of the section, it may reproduce variations in the training data at the local patch level.
+At a larger scale, this effect was pronounced in the form of stripe patterns, for example in the primary motor cortex ROI (Fig.˜5).
+These patterns typically coincide with inhomogeneities in the 3D-PLI transmittance maps.
+During the mounting process of the 60 µm thick sections in 20% glycerol solution on the object slide the gray matter is more prone to swelling than the more compact white matter, leading to tissue expansion, mechanical stress and tape flutter especially in the cortex due to its higher flexibility.
+The free-floating tissue between object slide and coverslip shapes waves of up to 30 µm in amplitude.
+These waves affect the focus level and the light incidence on the tissue, hence the effective birefringence, modifying the polarization state of the transmitted light.
+This caused the observed stripe patterns in the transmittance maps and very likely contributed to the staining inhomogeneities.
 Cell body membranes and blood vessel walls pose hard edges in the cortex, resulting in diffraction patterns even brighter than the background field.
 Varying of the focus level of the LMP-1 microscope allows optical scanning of the Poisson spots induced by diffraction.
-Cell bodies are highlighted with higher contrast by shifting and lowering the focal plane to the top and the bottom of the tissue by approximately $\pm$30 µm (Fig.
-11).
-This means that the ability to predict cells from 3D-PLI measurements is sensitive to the focus level of the microscope, and it is difficult to generalize to unseen focus levels that were not part of the model training.
-During the mounting process of the 60 µm thick sections in 20% glycerol solution on the object slide the gray matter is more prone to swelling than the more compact white matter, leading to tissue expansion, mechanical stress and tape flutter especially in the cortex due to its higher flexibility.
-The free-floating tissue between object slide and coverslip shapes waves of up to 30 µm in amplitude which affects the focus level and the light incidence on the tissue, hence the effective birefringence, modifying the polarization state of the transmitted light.
-Therefore, the stripe-like inhomogeneities in the virtual staining can be explained by such variations in focus levels within the same section.
-An effective countermeasure against waves in the tissue is to weight the open coverslips after embedding overnight.
-In addition, the acquisition of multiple focus levels for each section could potentially provide missing information.
+Cell bodies are highlighted with higher contrast by shifting and lowering the focal plane to the top and the bottom of the tissue by approximately $\pm$30 µm (Fig.˜11).
+The ability to predict cells from 3D-PLI measurements was shown to be sensitive to variations in the focus level.
+An effective countermeasure against waves in the tissue, to maintain a constant level across the entire section, is to weight the open coverslips after embedding overnight.
+In addition, the acquisition of multiple levels for each section could potentially provide missing information.
 For the given dataset, however, these potential solutions were not applicable, and measurements could not be easily repeated.
-In addition, the acquisition of multiple focus levels would significantly increase the measurement effort.
+Since shifting the focus toward the tissue surface degrades optical resolution, a center focus level is used by default in 3D-PLI for an optimal representation of nerve fibers.
+The acquisition of multiple focus levels of $\pm$30 µm to increase the visibility of cell bodies would significantly increase the measurement effort.
 Nevertheless, it will be an important consideration for future investigations to include multiple focus levels in each measurement to improve cell extraction.
+It must be noted that shrinking and swelling of the tissue also affects the measured areas of individual 2D segmented cell bodies.
+Tissue deformation in histology is complex, involving anisotropic effects and differences between cellular and extracellular structures dorph-petersen2001; west2013, which cannot be fully captured by the global shrinkage factor estimated in Sec.˜2.1.3.
+At the same time, there is only sparse data available that quantifies such effects, and the precise conditions of histological processing differ between each other, and it can be assumed that such differences influence the specific shrinkage values.
+In paraffin-embedded human sections, a mean difference of 9% between gray and white matter shrinkage has been reported kretschmann1982.
+As frozen sections generally show less in-plane deformation than paraffin sections (west2013), we treat this value as an upper bound for our acqusition.
+Since gray matter contains relatively more cell bodies than white matter, similar shrinkage of gray and white matter suggests that shrinkage of cell bodies and neuropil should also not differ significantly amunts1999.
+By performing non-linear co-registration of Cresyl violet to 3D-PLI and applying a subsequent global correction of cell sizes to account for swelling effects from MRI to 3D-PLI, we assume that major shrinkage effects are compensated.
+However, residual local variations or relative slice-to-slice differences may persist and will require a more systematic approach in future work.
 While most of the larger cells (>100 µm² in-plane size) could be localized in 3D-PLI parameter maps, smaller cells were often missed or misplaced in the predictions.
 This is expected as smaller cell structures are often dominated by a stronger signal of intersecting nerve fibers or overlay each other due to the high section thickness of 60 µm required for 3D-PLI.
-The predictions also showed a stronger arrangement of cells into cortical columns than in the target images.
+In our experiments, we observed that stronger birefringence, corresponding to higher retardation values, resulted in reduced capability of models to reconstruct cell instances.
+Since birefringence increases with myelin density and homogeneity, predictions in regions with densely packed myelinated nerve fibers, such as white matter, are thus less reliably captured by the virtual stainings.
+In white matter, detecting cell bodies may even be infeasible, as the predominant oligodendrocytes form the myelin sheath and are therefore indistinguishable from it.
+In gray matter, this limitation may explain the stronger arrangement of cells into cortical columns in the predictions compared to the target images in Fig.˜8, as the strong signal of myelinated radial fibers can obscure underlying cell instances.
 The presence or absence of this arrangement is one of the criteria used to identify cortical areas, which might impair downstream interpretation of the predicted architecture.
-It may be attributed to the pronounced signal of radial fibers in 3D-PLI, which are used by the virtual staining model to infer potential cell locations.
-Further investigations are necessary to understand this effect.
-Our ablation experiments in Sec.
-3.2, Sec.
-3.3 and Sec.
-3.4 support the design decisions of the proposed virtual staining method.
-They showed an overall good reconstruction and localization of individual cell bodies.
-But, of course, the virtual stainings can not yet fully replace real measurements.
-They occasionally include staining inhomogeneities, miss smaller cells, or introduce implausible cytoarchitectonic features, such as artificial columnar arrangements of cells.
-In contrast, extracting GLI profiles from the predictions demonstrated that the laminar cell organization can be reconstructed quite well.
-However, further investigations with larger datasets spanning multiple brain areas are necessary to thoroughly assess their reliability.
-The proposed methodology might be further improved by extending the amount of training sections, brain regions and focus levels in 3D-PLI acquisition.
-We also expect that the expressiveness of the Gram matrix representation could be enhanced by replacing the VGG encoder, pre-trained on ImageNet, with a domain-specific model, which is powerful on microscopy data.
-This is technically motivated by the observation that higher-level features in ImageNet-pretrained networks, such as VGG, are optimized for object semantics in natural images (e.g., faces, animals, vehicles), which are irrelevant in the context of fine-grained texture characteristic of microscopy images Stuckner et al., (2022).
-In contrast, microscopy-specific encoders are better suited to capture such low- to mid-level textural features and have been shown to improve clustering Oberstrass et al., (2024), segmentation Stuckner et al., (2022), and the evaluation of generative models Kropp et al., (2024) over those trained on ImageNet.
-Therefore, computation of Gram matrices using domain-specific encoders Spitzer et al., (2017); Schiffer et al., 2021b ; Schiffer et al., 2021a or emerging foundation models for histology Tran et al., (2024) might further enhance representation of cytoarchitectural characteristics.
+As shown by the differences in F1 values in Tab.˜5, cells were reconstructed more accurately in isocortical areas (motor cortex and temporal cortex) than in the hippocampus or the subcortical nuclei (putamen and globus pallidus).
+The lower F1 in the putamen compared to the motor and temporal areas can be attributed to its neuronal composition and the fundamentally different spatial arrangement of cell bodies in cortical versus subcortical regions, as well as their relation to fiber bundles.
+The putamen is composed of medium-sized spiny neurons and relatively small interneurons arranged around and between bundles of myelinated fibers, which course through the putamen and give rise to its characteristic striated appearance heilbronner2025.
+In this region, two challenges reduce predictability: small interneurons are difficult to detect with our approach, and fiber bundles may intersect or obscure cell bodies, making them harder to reconstruct.
+Moreover, the cell bodies of spiny neurons can resemble fiber bundles oriented perpendicular to the section plane, appearing as dark patches in the transmittance image.
+The globus pallidus contains relatively large but sparsely distributed neurons, which are embedded in a dense matrix of multidirectionally oriented myelinated fibers heilbronner2025.
+Here, predictability is reduced by the number and complex spatial organization of intersecting myelinated fiber bundles and blood vessels.
+The low F1 value in patches extracted from the CA1-CA3 regions of the hippocampus is more difficult to explain, though it may result from the fact that they cover both the pyramidal layer, with its numerous and relatively large pyramidal cells, and the radiatum layer, which contains sparsely packed and relatively small interneurons zhao2025, which are difficult to detect with our method.
+Our Gram+Reg model produced larger cells at expected positions and generated images with a plausible appearance, making virtual stainings useful for cross-modality registration - at present the most relevant application.
+They also allowed the application of cytoarchitectonic tools such as cell segmentation or computation of GLI profiles for cortical layer characterization on 3D-PLI images.
+As a future perspective, the ability to detect cell bodies in 3D-PLI may enhance the computation of 3D fiber orientations within gray matter, which is a prerequisite for fiber tractography.
+This could be achieved by improving the estimation of myelination, for example through the identification of voxels dominated by cell bodies, and potentially also support the localization of axon terminals.
+However, virtual stainings occasionally contained artifacts, including staining inhomogeneities, omission of smaller cells, or the introduction of implausible cytoarchitectonic features.
+These issues may reflect biases in the training data, model-related artifacts, or missing cellular signatures in the 3D-PLI parameter maps.
+The virtual staining allowed the identification by a neuroanatomist (N.P.-G.) of the borders between the CA1, CA2, and CA3 regions of the hippocampus, the border between the primary motor and primary somatosensory areas, or the border between the retrosplenial cortex and cingulate area 23.
+However, it was not possible to identify the border between the core and lateral belt auditory areas, because it is characterized by differences in the packing density of small pyramids hackett2001, which currently cannot be reliably predicted.
+Therefore, we do not yet consider the method sufficiently robust for reliable cytoarchitectonic brain area mapping.
+Nevertheless, we are convinced that expanding the number of training sections, brain regions, and focus levels across brains and species can improve model performance and generalizability.
+While such expansion would require re-training, each additional sample would teach the model novel architectural patterns, improving its robustness across domains.
 5 Conclusions
-Motivated by previous observations that larger cells are encoded in 3D-PLI parameter maps alongside fiber orientations Zeineh et al., (2017), we introduced a deep learning model for transforming 3D-PLI maps into virtual Cresyl violet cell body stainings.
+Motivated by previous observations that larger cells are encoded in 3D-PLI parameter maps alongside fiber orientations zeineh2017, we introduced a deep learning model for transforming 3D-PLI maps into virtual Cresyl violet cell body stainings.
 This approach enables joint visualization of fiber tracts and cell bodies in the same tissue.
-Compared to real post-staining, the model offers a scalable alternative that avoids manual labor, limited only by its generalization to unseen architectural patterns.
-Validation against post-stained test sections confirms that our virtual stainings reliably localize most of the larger cell bodies (>100 µm² in-plane size).
+Compared to real post-staining, the model may offer a scalable alternative that avoids manual labor.
 A central contribution of our approach is the integration of an online registration head during training.
-This component eliminates the need for explicit multimodal registration, which is commonly required in virtual staining pipelines Rivenson et al., (2020); de Haan et al., (2021); Yang et al., (2022).
+This component eliminates the need for explicit, pixel-accurate multimodal registration, which is commonly required in virtual staining pipelines rivenson2020; dehaan2021; yang2022.
 It is a simple but highly efficient add-on that can be combined with various loss formulations, leveraging model-estimated landmarks, to continuously refine the alignment over time.
-The method enables accurate prediction of cell bodies from 3D-PLI and successful adaptation to the appearance of real Cresyl violet stainings.
+The developed method enables localization of most of the larger cell bodies in gray matter (>100 µm² in-plane size) from 3D-PLI and a successful adaptation to the appearance of real Cresyl violet stainings.
 As such, it expands the usability of 3D-PLI in large-scale data settings, allowing virtual staining at scale.
 While such synthetic data cannot and should not replace real histological measurements, it offers promising opportunities in downstream analysis.
-Potential applications include cross-modal image registration to align real Cresyl violet and 3D-PLI images, disentangling signal sources to identify voxels dominated by cell bodies (e.g., for improved tractography), and missing data imputation in serial section stacks.
+Although we see the present method not yet robust enough for cytoarchitectonic mapping, potential applications include cross-modal image registration to align real Cresyl violet and 3D-PLI, performing cell segmentation in 3D-PLI images, or missing data imputation in serial section stacks.
 Especially in interleaved modalities, this could enable the reconstruction of complete datasets.
-Of course, such applications require careful quality control and distinction between synthetic and real measurements.
+Of course, such applications require careful quality control, a clear separation and demarcation of synthetic data from real measurements, and careful interpretation of derived results.
 The outcomes of this study lay the groundwork for prospective investigations focused on enhancing 3D-PLI analysis, particularly through the exploration of dedicated cell detection techniques to directly extract cell body instances from 3D-PLI data.
@@
 Code for the online registration  [^5] https://jugit.fz-juelich.de/inm-1/bda/software/data_processing/ffreg, data augmentations for 3D-PLI images  [^6] https://jugit.fz-juelich.de/inm-1/bda/software/data_processing/pli-transforms, visualization methods for 3D-PLI modalities  [^7] https://jugit.fz-juelich.de/inm-1/bda/software/data_processing/pli-styles, as well as additional dependencies  [^8] https://jugit.fz-juelich.de/inm-1/bda/software are hosted on our external GitLab server.
-The ROIs employed for the training and testing of the models in this study, along with a selection of model predictions, are available in our central institutional repository Oberstrass et al., (2025).
+ROIs employed for the training and testing of the models in this study, along with a selection of model predictions, are available in our central institutional repository oberstrass2025dataset.
+In addition, the repository includes whole-slide predictions for test section 559 along with corresponding 3D-PLI modalities.
 Author Contributions
@@
 Timo Dickscheid: Conceptualization, Methodology, Supervision, Writing - original draft, Writing - Review & Editing, Project administration, Funding acquisition.
-Conflict of Interest Statement
+Declaration of Competing Interests
 The authors declare that the research was conducted in the absence of any commercial or financial relationships that could be construed as a potential conflict of interest.
@@
 Appendix A Appendix
-Figure 11: Virtual staining quality varies with the focus level of the LMP-1 due to differences in the captured diffraction patterns.
-Example images show the vervet entorhinal cortex with zoom-ins to the architecture of cortical layers II-III.
-3D-PLI inputs are represented by the transmittance $I_{T}$, retardation $\sin\delta$ (scaled with gamma correction) and fiber orientation in HSV color space (hue: direction $\varphi$; saturation, brightness: retardation $\sin\delta$).
-The same Gram+Reg model is used to predict virtual stainings from a focus level manually adjusted to the tissue center (+0 µm) and 30 µm above and below.
-A lower focus level of -30 µm achieves the highest consistency in staining intensity and produces the most cells.
-Higher focus levels introduce an increasing amount of staining artifacts.
+Table 5: Specific quantitative results per test ROI, corresponding to averages reported in Tab.˜2.
+The deviation is reported as standard error over four independent trainings with different random seeds.
+Arrows indicate the direction of better performance ($\uparrow$ higher is better, $\downarrow$ lower is better).
+Best scores per ROI in bold.
+ROI Method MI $\uparrow$ RMSE $\downarrow$ SSIM $\uparrow$ F1 $\uparrow$
+Motor cortex Gram 0.061 $\pm$ 0.013 33.9 $\pm$ 1.0 0.306 $\pm$ 0.024 24.8 $\pm$ 3.5
+Gram+Reg 0.168 $\pm$ 0.003 29.2 $\pm$ 0.6 0.446 $\pm$ 0.002 42.9 $\pm$ 0.3
+GAN 0.117 $\pm$ 0.016 34.7 $\pm$ 1.1 0.268 $\pm$ 0.010 16.8 $\pm$ 1.0
+GAN+Reg 0.165 $\pm$ 0.015 29.6 $\pm$ 1.0 0.394 $\pm$ 0.013 36.3 $\pm$ 2.1
+Hippocampus Gram 0.184 $\pm$ 0.015 34.5 $\pm$ 0.9 0.340 $\pm$ 0.013 19.4 $\pm$ 2.3
+Gram+Reg 0.303 $\pm$ 0.003 31.1 $\pm$ 0.5 0.448 $\pm$ 0.001 31.4 $\pm$ 0.4
+GAN 0.078 $\pm$ 0.016 44.3 $\pm$ 3.1 0.262 $\pm$ 0.016 6.9 $\pm$ 1.3
+GAN+Reg 0.151 $\pm$ 0.036 32.1 $\pm$ 0.6 0.373 $\pm$ 0.027 24.9 $\pm$ 1.9
+Temporal cortex Gram 0.077 $\pm$ 0.018 33.9 $\pm$ 0.8 0.319 $\pm$ 0.040 31.7 $\pm$ 4.5
+Gram+Reg 0.236 $\pm$ 0.003 29.1 $\pm$ 0.4 0.497 $\pm$ 0.004 52.7 $\pm$ 0.4
+GAN 0.051 $\pm$ 0.006 43.1 $\pm$ 5.4 0.238 $\pm$ 0.019 18.3 $\pm$ 2.7
+GAN+Reg 0.192 $\pm$ 0.015 30.1 $\pm$ 1.5 0.460 $\pm$ 0.017 49.6 $\pm$ 1.3
+Subcortical nuclei Gram 0.122 $\pm$ 0.007 32.0 $\pm$ 0.6 0.290 $\pm$ 0.016 15.7 $\pm$ 2.2
+Gram+Reg 0.190 $\pm$ 0.004 29.2 $\pm$ 0.4 0.388 $\pm$ 0.001 29.4 $\pm$ 0.5
+GAN 0.138 $\pm$ 0.028 33.0 $\pm$ 0.8 0.245 $\pm$ 0.013 11.4 $\pm$ 0.9
+GAN+Reg 0.132 $\pm$ 0.014 29.9 $\pm$ 0.7 0.330 $\pm$ 0.014 20.9 $\pm$ 1.6